### v1

# FlexGuard: A Design Space for On-Body Feedback for Injury Prevention in Strength Training

Abstract.
Strength training carries risk of injury when exercises are performed without supervision. While haptics research has advanced, there remains a gap in how to integrate on-body feedback into intelligent wearables. Developing such a design space requires experiencing feedback in context, yet obtaining functional systems is costly. By addressing these challenges, we introduce FlexGuard, a design space for on-body feedback to support injury prevention in strength training. The design space was derived from nine co-design workshops, where novice trainees and expert trainers DIY’d low-fidelity on-body feedback systems, tried them immediately, and surfaced needs and challenges encountered in real exercising contexts. We then evaluated the space through speed dating, using storyboards to cover the design dimensions. We followed up with workshops to further validate selected dimensions in practice through a proof-of-concept wearable system prototype. Our findings extend the design space for sports and fitness wearables in the context of strength training.
1. Introduction
Strength training is a widely practiced method for improving strength, mobility, and overall health. However, errors such as joint misalignment or overextension can increase the risk of injury, especially for novice practitioners (Fisher, 2025). Personal trainers help mitigate these risks by providing tailored feedback, but such supervision is costly and depends on others’ availability, making it difficult to scale. Advances in haptics now enable the emulation of sensations such as taps, pressure, and resistance. This creates an opportunity for intelligent wearable systems that can deliver embodied feedback to support safe and effective training. What remains missing, however, is the design knowledge needed to integrate on-body feedback into such systems.
Designing on-body feedback systems is challenging because users have to physically experience feedback during strength training (Tennent et al., 2020), and obtaining functional systems is costly (Caserman et al., 2021). Investing heavily in a prototype to serve as a technology probe (Hutchinson et al., 2003) is often impractical in early-stage exploration. Moreover, any probing system must be representative of the diverse on-body feedback mechanisms that have been studied and developed. Together, these challenges hinder systematic investigation of the design space.
By addressing these challenges, we developed FlexGuard, a design space for on-body feedback in strength training injury prevention. We began by surveying the landscape of feedback modalities, including vibrotactile (Huang et al., 2025), pneumatic (Delazio et al., 2018; Pardomuan et al., 2024), and other approaches (Kwon et al., 2022; Lin et al., 2025b), and grouped them into two categories: extrinsic, which require active actuation (e.g., vibration, air pressure), and intrinsic, which rely on material properties (e.g., orthoses, braces) to provide passive feedback once configured. We then prepared technology probes representing these categories, along with supporting props, and conducted nine co-design workshops in our institution’s gym. Each session paired a novice trainee with an expert trainer, who collaboratively DIY’d personalized low-fidelity on-body feedback systems for five representative exercises and tested them immediately in practice. Through thematic analysis of session recordings and worksheets, we distilled the needs and challenges that informed the derivation of FlexGuard.
FlexGuard is organized around two complementary dimensions: Sensing & Triggering and Feedback Delivery. Sensing & Triggering specifies which indicators should be monitored (e.g., joint angle, movement dynamics, motion path, muscle activation, inter-limb coordination), how triggering policies are defined (e.g., expert-based or personalized), and where monitoring should occur (e.g., target, stabilizing, moving, or compensation regions). Feedback Delivery describes how feedback is provided once triggered, including the body zones where different feedback types are applied, the intensities of feedback, its directional strategies, and its timing across movement phases. Together, these dimensions define how on-body feedback can be timed, targeted, and configured.
We ran three speed dating workshops with seven trainees with limited strength training experience. Participants first performed exercises with the provided equipment in our lab to ground their reflections, and then discussed a series of storyboards that varied Sensing & Triggering and Feedback Delivery, spanning the dimensions of FlexGuard. The results revealed that participants interpreted feedback modalities with distinct semantics, emphasized the importance of feedback timing and intensity, and highlighted adaptability as critical across phases of movement, exercises, and user states. They also reflected on learning and safety, noting concerns that excessive support could reduce learning and that locking mechanisms could pose risks. Overall, these findings validated the proposed dimensions of FlexGuard while also pointing to areas for refinement.
We further validated the dimensions of FlexGuard in practice by developing a proof-of-concept (PoC) wearable system. Motivated by the speed dating findings, the PoC focused specifically on adaptability. Given the underexplored nature of adaptive intrinsic feedback, we narrowed the design to intrinsic mechanisms, implementing a strap-based suit tested on two exercises that together cover abduction–adduction and flexion–extension degrees of freedom of the shoulder. A camera-based system detected deviations from target posture and signaled strap adjustments, which were manually enacted in this PoC to simulate an automated system. This setup allowed participants to experience phase-specific variations in feedback direction and intensity, providing finer-grained validation of how the design space principles translate into embodied training contexts at the phase level.
Our contributions are as follows:
We introduce FlexGuard, a design space for on-body feedback in strength training, grounded in nine co-design workshops with trainers and trainees.
We articulate dimensions that characterize how systems can sense and deliver feedback to reduce injury risk across the exercises studied in the workshops.
We present two complementary evaluations: a speed dating workshop and a PoC wearable system, together demonstrating how FlexGuard can guide the design of adaptable on-body feedback systems.
2. Related Work
Our research builds upon prior work in four key areas: Injury Prevention in Strength Training, Sensing and Feedback Systems for Physical Training, Advances in Materials and Wearable Technologies, Intelligent Systems in Movement and Skill Learning.
2.1. Injury Prevention in Strength Training
Strength training carries inherent risks of injury, particularly when exercises involve high external loads that stress the musculoskeletal system. Body regions that both bear the load and move dynamically, such as the lower back and shoulders, are especially prone to injuries like strains, sprains, and impingements (Hedrick and Wada, 2008; Barnes et al., 2021). While such injuries are common, they are often avoidable. Research indicates that a large portion of lower back injuries are linked to improper lifting form, excessive loads, and overtraining—all factors that can be mitigated through proper technique and training practices (Alqarni, 2019). Faulty alignment and compensatory movement patterns introduce abnormal stress across joints and soft tissues, contributing not only to acute injury but also to long-term issues. This risk is especially pronounced in novice trainees, whose motor control systems may not yet be developed enough to coordinate complex compound movements effectively (Aslam et al., 2025). Thus, consistent attention to motor learning is fundamental for safe and effective strength training (Latella et al., 2019).
This is why working with a personal trainer is widely recommended, particularly for beginners. Professional supervision ensures that trainees are taught correct movement techniques, are properly spotted to maximize repetitions safely, and are guided through appropriate loading progressions. When properly coached, strength training is a safe activity with a low incidence of injury relative to other sports (Fisher et al., 2023). In contrast, unsupervised training is associated with higher risks due to inconsistent technique and overly aggressive load selection (Faigenbaum and Myer, 2010; Gabbett, 2020). While technologies exist to support self-coaching (Snouzy, 2025), human coaches provide multifaceted guidance that extends beyond form correction and motor habit development to include enforcing safety practices such as spotting, load monitoring, and early detection of execution breakdowns (Fisher et al., 2023).
In addition to coaching, trainees frequently use supportive wearable gear to enhance injury prevention. Equipment such as weightlifting belts, wrist wraps, knee sleeves, and elbow sleeves is commonly used in gyms to help maintain alignment and distribute stress during heavy lifts. Weightlifting belts, for instance, are used to increase intra-abdominal pressure and reduce spinal loading during squats and deadlifts (Herbaut and Tuloup, 2025). Knee sleeves improve proprioception during deep knee flexion movements (Machek et al., 2021). Users can also define and produce custom wearables for their specific needs (Xu et al., 2025). For example, tailoring their own knee patch or squat suit to support heavy lifting (Latella et al., 2024).
FlexGuard builds on this trajectory by extending injury prevention beyond passive support and human supervision. Leveraging haptics for on-body sensation, FlexGuard advances toward adaptive systems that actively sense and respond to cues and deviations during exercise, providing intuitive feedback to promote safer training.
2.2. Sensing and Feedback Systems for Physical Training
Designing wearables requires careful consideration of how the body perceives and responds to interaction, emphasizing not just function but felt experience (Tennent et al., 2020); thus, there are two key components: sensing and feedback methods on the body. Sensing technologies have enabled researchers and designers to monitor a wide range of body states during physical training. Researchers have explored the use of sensors to monitor muscle activity (Zhu et al., 2022; Esposito et al., 2018; Rudolph et al., 2022). Systems like RecoFit (Morris et al., 2014) and LiftSmart (Kowsar et al., 2019) use wearable sensors to monitor and assess users’ exercise performance. Other systems rely on pressure or tactile sensors to provide posture-related sensing, such as GymSoles, which uses insole-based pressure readings to visualize a user’s center of pressure during lifts (Elvitigala et al., 2019). FormFit (Rao et al., 2024) employs computer vision for automatic pose correction, while Zishi (Wang et al., 2016) combines smart textiles and motion sensing to monitor posture during rehabilitation. More integrated systems like E-Orthosis (Yang et al., 2023) embed sensing directly into orthotic devices using textile electrodes, enabling real-time feedback and data logging in rehabilitation and sports contexts. Researchers have also studied textile-based knee strain sensors for real-time monitoring during workouts (Aliyana et al., 2024), as well as developable shoulder sensors for precise movement tracking (Evke et al., 2019). Commercial platforms such as Catapult (Catapult, 2025) offer elite-level sports analytics, although they often remain inaccessible or inflexible for everyday users.
A parallel stream of work focuses on feedback modalities for physical training. Researchers have extensively explored visual augmentation through projection and AR/VR (Yu et al., 2024), including systems such as BodyLights (Turmo Vidal et al., 2020), augmented instructional videos (Semeraro and Turmo Vidal, 2022), real-time overlays (Keelawat and Suzuki, 2024; Ma et al., 2024), and spatial visualizations in FlowAR (Jo et al., 2023a), all of which guide users with contextual cues. Some systems, like Weight-Mate (Paay et al., 2020), combine visual and auditory feedback to help users maintain alignment. Haptic feedback systems have also gained attention, including vibration-based alerts for physical activities (Spelmezan et al., 2009; Kosmalla et al., 2016), implemented in TrainerTap (Jo et al., 2023b) VibraForge (Huang et al., 2025) and Subtletee (Woźniak et al., 2020). Such systems may be particularly beneficial for individuals with visual impairments when practicing yoga (Islam et al., 2022). Han et al. (Han et al., 2023) explore 3D-printed tactor geometries for generating diverse haptic sensations. EMS-based stimulation for movement correction is another promising direction; for instance, Teslasuit (Caserman et al., 2021) exemplifies a production-grade electro muscle stimulation (EMS) and transcutaneous electrical nerve stimulation (TENS) platform capable of training full-body motion. Other systems explore novel modalities, including skin-stretching for proprioceptive illusions (Lin et al., 2024b), pneumatic water-jet feedback (Zhang et al., 2024), localized pressure for knee joint stabilization (Park et al., 2020), and torque-based force guidance for arm motion (Kim and Asbeck, 2022).
Researchers have studied broader challenges in the context of sports technologies (Turmo Vidal et al., 2021; Elvitigala et al., 2024), highlighting the need for systems that can adapt to individual physical contexts and training goals. Personalization is especially emphasized in recent work on sports wearables (Stawarz et al., 2025), while others call for more interpretability and objective relevance in how feedback is delivered (Tajadura-Jimenez et al., 2022).
These works show the breadth of sensing and feedback modalities available, but most rely on predefined mappings and offer limited user configurability. Even state-of-the-art products like the Teslasuit (Caserman et al., 2021) require heavy reliance on expert setup. FlexGuard addresses this gap by enabling end users to configure, time, and adjust feedback delivery interactively, while maintaining safety by anchoring its logic in expert knowledge.
2.3. Advances in Materials and Wearable Technologies
Recent developments in soft robotics, programmable materials, and digital fabrication have expanded the design space for wearable systems. Pneumatic actuation systems like PneumAct (Günther et al., 2019), VabricBeads (Pardomuan et al., 2024), Force Jacket (Delazio et al., 2018), and Frozen Suit (Al Maimani and Roudaut, 2017) use air or liquid pressure to generate movement, resistance, or stiffness changes in the body. milliMorph (Lu et al., 2019) pushes this further by introducing millimeter-scale fluidic chambers for fine-grained, fluid-driven shape change. PneuMa (Saini et al., 2024) introduces pneumatic bodily extensions to support daily movement. Actuation mechanisms inspired by biological forms, such as artificial muscles (Li et al., 2017), offer compact and powerful actuation possibilities. Researchers have also investigated collaborative approaches for designing wearables with on-body feedback, such as vibrotactile patterns in remote scenarios where designers are not co-located (Wittchen et al., 2025).
Programmable materials with bistable or multistable properties allow for richer forms of mechanical feedback and passive guidance (Alu et al., 2025). Works such as Kirigami Haptic Swatches (Chang et al., 2020), FlexTure (Johnson et al., 2023), Waxpaper Actuator (Wu et al., 2024), and ConeAct (Lin et al., 2024a) explore structural snapping transitions, geometric locking, and reactive haptics. Snap-through mechanisms, like those in SnapInflatables (Yang et al., 2024), enable rapid shape changes with localized feedback. Springlets (Hamdan et al., 2019) introduces shape-memory alloy springs that simulate on-skin sensations such as touching, pinching, pulling, etc.
Wearable fabric-based sensing and actuation is also advancing rapidly (Li and Luo, 2024). Luo et al. demonstrate textile-based control for skill transfer tasks like piano playing (Luo et al., 2024), while ReKnit-Care (Hu et al., 2025) combines knitted sensors with EMS to provide fine-grained hand feedback for rehabilitation. Lin et al. (Lin et al., 2025a) introduce passive microstructures to switch between states while worn to change perceived physical environment. Tessmer et al. designed a spacesuit with tunable compression (Tessmer et al., 2022). FlexGuard draws from this material ecosystem, integrating advances in actuation and textiles to explore how adaptive, on-body feedback can support safe and personalized strength training.
2.4. Intelligent Systems in Movement and Skill Learning
Interactive Machine Learning (IML) refers to systems that allow end-users to guide, correct, or shape the learning process through iterative interaction. Unlike traditional supervised ML, IML invites user-in-the-loop contributions during model refinement (Amershi et al., 2014). Hartmann et al. (Hartmann et al., 2007) emphasize the value of interactivity in enabling more intuitive and user-aligned learning. These systems combine sensor data, user feedback, and pattern recognition to empower non-expert users to build and refine interactive logic.
In movement learning and demonstration-based systems, researchers have applied IML techniques to encode and reuse human motion, which often involves tacit knowledge that is difficult to articulate explicitly (Gillies, 2019). MotionMA (Velloso et al., 2013) enables users to demonstrate movement patterns that the system can model and compare in real time. Mimic (Mahadevan et al., 2022) applies similar principles in teleoperation, letting users record and re-trigger physical behaviors. Other approaches like danceON (Payne et al., 2021) and Rapido (Leiva et al., 2021) focus on embodied programming, through specification or demonstration, for creative and spatial applications. Vidal et al. propose a framework and infrastructure for applying IML in physiotherapy to support both in-clinic and out-of-clinic patients (Turmo Vidal et al., 2025). Haptic Scores (Escudero, 2024) attempts to map detected movements to embodied haptics for dance with EMS feedback. In sport-specific contexts, Weng et al. (Weng et al., 2025) decompose movement into feedback-friendly units to integrate AI for teaching basketball shooting skills. While recent advancements in large language models (LLMs) have significantly expanded the capabilities of intelligent systems (Mahadevan et al., 2024), designing intuitive programming environments that effectively integrate LLMs remains a challenge (Pu et al., 2025). FlexGuard builds on these directions by embedding interactive intelligence directly into the wearable interface, enabling trainees to shape feedback for safer and more personalized strength training.
3. Landscape of Feedback Modalities
Wearable systems have explored a wide range of on-body feedback modalities to support physical interaction and learning. Our review of over 50 papers from HCI research, complemented by examples from gym practices and selected work in soft robotics, identifies recurring strategies including vibration, electrical stimulation, pneumatic actuation, thermal cues, material-driven mechanisms, and structural supports. Table 1 summarizes representative work across categories.
| Modality | Examples | Description |
| Vibration | TrainerTap (Jo et al., 2023b), VibraForge (Huang et al., 2025), HapticHead (Kaul and Rohs, 2017), Zishi (Wang et al., 2016) | Pulses or patterns on the skin. |
| Electrical (EMS / TENS) | Teslasuit (Caserman et al., 2021), Smartwatch EMS (Takahashi et al., 2024), ReKnit-Care (Hu et al., 2025), Haptic Scores (Escudero, 2024) | Electrical stimulation of muscles or nerves. |
| Pneumatic | VabricBeads (Pardomuan et al., 2024), PneumAct (Günther et al., 2019), Force Jacket (Delazio et al., 2018), FrozenSuit (Al Maimani and Roudaut, 2017), Bellowband (Young et al., 2019), JetUnit (Zhang et al., 2024) | Air or fluid pressure for force or resistance. |
| Thermal | ThermoVR (Peiris et al., 2017), Augmented Breathing (Brooks et al., 2024) | Heating or cooling sensations. |
| Material-Driven | Springlets (Hamdan et al., 2019), ArmDeformation (Lin et al., 2024b), ImpactVest (Tsai et al., 2022), OmniFiber (Kilic Afsar et al., 2021) | Mechanical actuation or material deformation. |
| Structural Supports | Bistable Orthoses (Lin et al., 2025b), Stiffening Garment (Kwon et al., 2022), SkinMorph (Kao et al., 2018), E-Orthosis (Yang et al., 2023), gym gear (e.g., knee sleeves (Machek et al., 2021), squat suit (Latella et al., 2024), lifting belts (Herbaut and Tuloup, 2025)) | Passive supports for posture or training. |
Across these categories, we distinguish between extrinsic mechanisms (active modalities such as vibration, electrical, pneumatic, thermal, or material-driven actuation that deliver forces or sensations to the body) and intrinsic mechanisms (passive supports such as braces, sleeves, or garments that guide or resist movement through direct material contact). Since strength training demands both corrective cues and physical support, our subsequent sections of this paper focus on these two modalities through co-design workshops exploring their affordances in training contexts.
4. On-Body Feedback for Strength Training
To understand the needs and challenges of on-body feedback systems for strength training, we conducted a co-design workshop study with novice trainees, supported and informed by expert trainers. To achieve that goal, we have to prepare materials for the study including technology probe and worksheets, that will cover comprehensive space of strength training.
We focused on a set of upper-body strength exercises where errors could lead to injury, to scope the study. Although not full-body, the selected exercises encompass a broad range of movement demands in strength training, offering insights that can largely generalize to other body regions and exercises. The representative set included five exercises: dumbbell front hold, dumbbell incline fly, bench press, lateral raise, and Arnold press. Each exercise highlighted a distinct movement error risk, such as posture holding, range of motion (ROM), inter-limb balance, tempo, or complex transitions. Further details about injury risks are shown in Fig. 10.
4.1. Technology Probe Design for DIY Low-Fi Prototyping
To enable rapid DIY prototyping and immediate exploration of feedback sensations, we developed a technology probe (Hutchinson et al., 2003) designed for quick reconfiguration. The base system included adjustable shoulder braces and elbow sleeves, equipped with customizable feedback components such as inflatable air wedges on the shoulders, and rigid and stretchy straps on the elbows. The probe supported two distinct modes of on-body feedback:
4.1.1. Extrinsic Feedback
System-actuated and on-demand, simulated through the inflation of air wedges during exercise.
4.1.2. Intrinsic Feedback
Material-driven and continuously available, achieved through tension in rigid or stretchy straps that responded directly to movement.
Velcro interfaces allowed participants to quickly attach, reposition, or remove components as needed. To further support collaborative prototyping, we provided additional materials such as duct tape, resistance bands, and a mini-mannequin to facilitate communication between trainers and trainees. This setup enabled hands-on experimentation: participants configured feedback on their own bodies, performed representative movements, and reflected on the sensations in real time (Fig. 2).
4.2. Worksheet Design
We prepared a worksheet for each exercise. Each worksheet included body maps (Turmo Vidal et al., 2023) with figures of a person performing the exercise at representative moments, such as the starting position, mid-movement, and the maximum extension point. For the dumbbell front hold, an additional figure depicted an isometric hold at 15 seconds. These figures served as reference frames for participants to annotate, with the option of requesting extra sheets to sketch alternative designs or capture other parts of the movement cycle. As body maps are adaptable to diverse research contexts (Turmo Vidal et al., 2023), in our study they functioned both as a tool to guide the session and as an open-ended space for dialogue between trainer and trainee. They were designed for participants to sketch low-fidelity prototype designs, align on preferred configurations, and highlight specific body regions requiring attention. An example worksheet can be seen in Fig. 3.
4.3. Study Protocol
We conducted nine co-design sessions, each involving a matched trainer–trainee pair. Table 2 summarizes participant demographics, training experience, and notable athletic backgrounds.
| Session |
Other Sports & Fitness Background |
| S1 | E1 (48, M) | 2 yrs | 5+ | – | N1 (23, M) | |||||||||
| S2 | E2 (26, M) | 1.5 yrs | 5–10 | – | N2 (26, M) | |||||||||
| S3 | E3 (25, F) | 1 yr | 2 | – | N3 (33, F) | |||||||||
| S4 | E4 (28, M) | 1 yr | 2 | – | N4 (24, M) | |||||||||
| S5 | E5 (24, F) | 8 yrs | 5+ | • Licensed Zumba & Aerobics instructor • Experience teaching classes with handheld weights | N5 (26, F) | |||||||||
| S6 | E6 (23, F) | 9 yrs | 5+ | • Division I track/cross-country athlete • Provided instruction on form and spotting to peers and juniors | N6 (23, F) | |||||||||
| S7 | E7 (24, F) | 5-7 yrs | 20+ | • National-level rowing athlete • Instructor of rowing camps | N7 (33, F) | |||||||||
| S8 | E8 (22, M) | 8 yrs | 100+ | • Fitness instructor (3 years) • Track/cross-country athlete: PAC-12 Honor Roll, National U-20 champion | N8 (25, M) | |||||||||
| S9 | E9 (25, M) | 8 yrs | 30+ | • Gym trainer (5 months) • Trained peers in calisthenics club | N9 (31, M) |
All novice trainees (N1-N9) had less than three months of consistent strength training experience.
Participants were instructed to wear shirts with sleeves, as the technology probe was designed to be worn on the shoulders and arms. To ensure hygiene and minimize direct skin contact, we provided disposable elbow sleeves for participants to wear under the probe components.
Sessions took place in a private studio room at our institution’s gym and lasted approximately 90 minutes. To ensure participant comfort, especially during on-body probe configuration, each pair was matched by gender. All participants provided informed consent for video and audio recording. When participants used public gym equipment (e.g., for the bench press), continuous audio recording was maintained, and short video clips were recorded when appropriate, focusing only on consenting individuals. Throughout the session, participants were encouraged to think aloud and take breaks as needed. Each participant received $37.50 in compensation ($25 per hour) for their time and effort.
Each session followed the structure below:
4.3.1. Initial Interview
Each session began with two brief, separate audio-recorded interviews, one with the trainee and one with the trainer. Participants discussed their individual training backgrounds, any prior injuries from strength training, and their experience coaching or being coached. These interviews provided context for interpreting their design decisions during the rest of the session.
4.3.2. Worksheet Introduction
We introduced the study using worksheets to guide participants through the activities. We explained the project and highlighted three design goals to optimize:
DG1: Prevent injuries by providing feedback.
DG2: Deliver on-body sensations.
DG3: Personalize feedback to the individual needs of each trainee.
4.3.3. Probe Demonstration
Next, the research team introduced the technology probe. We first demonstrated the default configuration of the probe 2. Participants were then encouraged to experiment with customizing the setup to their own preferences, such as attaching straps in alternative positions (e.g., waist-to-shoulder). Both trainer and trainee wore the probe and explored three core feedback configurations:
Inflatable air wedges (extrinsic, active pressure feedback)
Rigid straps (intrinsic, non-elastic passive resistance)
Stretchy straps (intrinsic, elastic passive resistance)
This phase allowed participants to experience the tactile qualities of each configuration and develop an on-body sense of how feedback might support training. Since our focus was on personalizing feedback for trainees (DG3), trainers were not required to wear the probe beyond this stage.
4.3.4. Typical Training without Feedback
To establish a baseline and observe natural trainer–trainee dynamics, the trainee removed or loosened the probe, and the trainer guided them through a typical coaching session using minimal weight. Heavier weights were available to explore more challenging scenarios, but their use was optional, supervised by the trainer, and contingent on the trainee’s comfort and consent. Participants were encouraged to take breaks or drink water during the study. This phase helped surface real-world challenges, and implicit feedback strategies.
4.3.5. Co-Design Activity and Reflection
After the initial training, each trainer–trainee pair engaged in a collaborative co-design activity by sketching their ideas on the worksheet. They then configured the probe to reflect these design decisions, after which the trainee re-performed the exercise while wearing the configured probe. Participants reflected on how the feedback felt, discussed potential improvements, and recorded notes on the worksheet. If time permitted, they iterated on the design to better align with their intentions.
4.4. Data Analysis
We conducted a thematic analysis using reflexive approaches that combined both deductive and inductive coding (Braun and Clarke, 2019) of video and audio recordings as well as annotated worksheets. The first author developed initial codes informed by theoretical frameworks from prior review articles on sports wearable technologies (e.g., (Turmo Vidal et al., 2021; Seçkin et al., 2023)), focusing on themes such as inputs, outputs, and their interplay. After collaboratively reviewing two sessions with the fourth and seventh authors, including videos (Knoblauch et al., 2014), transcripts, and worksheets, the first author generated preliminary groupings, which were then shared with the second, third, fourth, and seventh authors to refine the emerging themes. These results were iteratively discussed and revised during team meetings with the broader author group and were also compared against more recent design space publications (e.g., (Stawarz et al., 2025)). The remaining recordings were reviewed by the fourth and seventh authors under the supervision of the first author. The resulting design space was extensively discussed and refined by the first, fourth, and seventh authors, with additional input from the last author as well as the rest of the team during meetings.
5. FlexGuard: A Design Space for On-Body Feedback for Injury Prevention in Strength Training
Drawing on insights from the co-design workshops, we introduce FlexGuard, a design space that models how on-body feedback can be sensed, triggered, and delivered to support safe strength training.
We structure this design space around five exercise phases: Initiation, Execution, Hold, Return, and Completion. The Execution and Return phases may be either concentric or eccentric, depending on the exercise. Dividing exercises into phases enables us to map specific design choices to context-dependent feedback opportunities.
FlexGuard comprises two overarching categories: Sensing & Triggering and Feedback Delivery. Throughout the paper, we use blue styling to denote dimensions of Sensing & Triggering, and purple styling to denote dimensions of Feedback Delivery.
5.1. Sensing & Triggering
Figure 4 summarizes the design space for Sensing & Triggering, spanning Trigger Metric, Trigger Policy, and Trigger Location. These dimensions define when and where the system should monitor to actuate feedback during strength training, as well as when to release it. Sensing must capture not only cues that warrant intervention but also signals indicating that feedback can be withdrawn once the user has corrected their form.
5.1.1. Trigger Metric
Participants identified several physical characteristics that could serve as reliable triggers for feedback. Across multiple sessions, joint angle was the most consistently mentioned marker, particularly when arms drifted out of alignment. As E5 noted, “your elbow keeps flaring out, that’s where it starts to go wrong.” Trainers also cautioned against end-range errors—“Again, you never want to lock your elbows” (S5). Others emphasized movement dynamics (e.g., acceleration or tempo) as indicators of excessive momentum rather than controlled lifting; during a lateral raise, E4 explained to N4, “if the dumbbells bounce at the top, that’s all shoulder stress instead of control.” In S1, participants highlighted inter-limb coordination (e.g., symmetric pacing/height in bench press) as a cue to synchronize both sides. In S8, feedback was tied to muscle activation, with the trainer remarking, “make sure you engage your chest muscle, otherwise you are not working out the muscles that you want.” Three sessions (S2, S7, S9) discussed motion path, noting that a shaky bar or wobbling dumbbell trajectory signals fatigue and breakdown of posture. For instance, when E2 noticed N2 developing a slight shake during the bench press, E2 intervened to help lift the bar.
5.1.2. Trigger Policy
Participants proposed rules for deciding when feedback should be actuated and when it should be withdrawn. Two broad categories emerged. expert-based policies draw on general training knowledge, such as recommended ROM or standard joint angles. For example, most trainers emphasized enforcing limits when elbows drift too far during a press, or constraining unsafe ROM at the extremes of an incline fly. In contrast, personalized policies adapt thresholds to the individual trainee, accounting for factors such as fatigue, muscle engagement, or body type. For instance, feedback could be tuned to a user’s tolerance for fatigue or to whether the intended muscle is activating effectively, both of which can vary significantly across body types. While all policies, including expert-based, can ultimately be personalized, we make this distinction to separate those with only slight variations across most body types from those that are highly subjective.
5.1.3. Trigger Location
Rather than naming specific muscles, we describe locations in terms of their functional roles to allow for greater generalizability. Four categories emerged, corresponding to Figure 4. The target region was a common focus, such as in the incline dumbbell fly: “If you started overcompensating and using your wrong muscle, it would like identify that muscle and you’d know you don’t want [that]” (S6). Trainers also warned about overuse of compensation regions, where non-target areas begin to take over the movement. The importance of stabilizing regions, particularly the upper back and core, was emphasized across sessions—“We always want to have them tight core. So tighten our core” (S5). Finally, participants pointed to moving regions (e.g., forearms and hands during an incline fly, or the wrists in a bench press) where feedback can guide trajectory directly. Although the muscle groups at these locations are not necessarily being worked, they act as reference points for maintaining correct movement.
5.2. Feedback Delivery
Feedback Delivery captures how corrective cues are conveyed to the user once a triggering condition is met as defined in Sensing & Triggering. We define this not only as the onset of feedback, such as vibration, resistance, or pressure, but also its offset. Participants emphasized that the withdrawal of feedback was equally meaningful, since the absence of cues could serve as confirmation that posture had been corrected or movement had returned to the desired path. In this way, Feedback Delivery encompasses both the activation and the release of feedback, as well as its modulation in intensity and timing.
5.2.1. Feedback Type
We distinguish between extrinsic feedback, delivered through active actuation, and intrinsic feedback, which emerges from material properties such as straps or elastic bands.
extrinsicfeedback can be tuned to produce different actuation patterns. Participants described preferences for continuous cues that guide movement throughout a repetition—“Imagine something helping draw a line for your hands to follow” (N2)—as well as discrete cues that act as one-time warnings when posture falters—“You shouldn’t nag them, just nudge once if they’re falling” (E1).
intrinsicfeedback depends on how materials constrain or guide motion. We distinguish between rigid and elastic responses. rigid feedback provides firm constraints, such as straps that enforce strict ROM limits: “I could feel when I pulled wrong because the strap caught me” (S5). In contrast, elastic feedback allows more flexibility, offering resistance that increases with deviation and guiding users back on track without completely halting movement. As E7 explained, “So a strap that is going to help them maintain the stability. When I’m pressing it, I don’t want to go up and down… if the strap helps us to do that, then it’s pretty good.”
5.2.2. Feedback Intensity
This dimension reflects how forceful the feedback feels, ranging from low to high. low intensity was favored for subtle prompts or early-stage correction. As E6 explained, “It should whisper, not shout.” In contrast, high intensity was valued for safety-critical moments, such as E4’s proposal for strong restricting feedback in the return phase of a lateral raise: “By the time they’re lowering, it better hit hard if they drop wrong.” Participants also emphasized progressive intensity—feedback that begins subtly and escalates if the error persists or worsens. This approach mirrors how a trainer might tolerate small deviations but intervene more forcefully over time as errors become more critical.
5.2.3. Feedback Latency
This dimension distinguishes between instant and delayed feedback, reflecting how quickly the system should respond once a deviation is detected. Immediacy is critical in situations involving high injury risk, where feedback must intervene before damage occurs. At the same time, participants recognized the value of delayed feedback that tolerates minor deviations and only escalates if errors persist. Such delayed feedback was seen as less disruptive and more supportive, as it felt less intimidating to users. Several also expressed preferences for progressive onset, where feedback begins subtly and intensifies if the deviation continues. Taken together, these insights suggest that trigger policies must balance the safety benefits of instant correction with the patience of delayed feedback that reduces cognitive overload.
5.2.4. Feedback Location
This dimension describes where on the body feedback is delivered. From the co-design workshops, we derived body zone maps that highlight regions associated with feedback (Figure 8). extrinsic feedback can be applied directly to the same muscle group identified by Sensing & Triggering, or redirected to a different site to signal correction. In contrast, intrinsic feedback is delivered at the site of motion itself, where resistance provides direct guidance. For example, constraining elbow flare or stabilizing the shoulders.
5.2.5. Feedback Direction
This dimension describes how corrective forces or cues are oriented relative to the body. We identified four strategies. opposing feedback resists unwanted momentum, such as adding friction in a lateral raise to counter swinging. guiding feedback nudges the user toward the correct path, for example assisting with rep completion or steering the arms back into the proper motion plane, as N6 described when correcting form during a dumbbell incline fly. attentional feedback provides localized cues without displacement, such as vibration in a dumbbell front hold to warn when the arms drop below a safe angle as suggested by N1. Finally, compressive feedback applies inward pressure, for instance stabilizing the core to maintain posture, which participants (e.g., E3, E4, E5) emphasized with their trainees, “always engage the core.”
5.3. Connecting Sensing & Feedback
The two dimensions of Sensing & Triggering and Feedback Delivery do not operate independently. Rather, they form a closed loop in which detected cues directly shape how feedback is delivered. Figure 10 illustrates this integration across five representative exercises surfaced during co-design workshops, with example approaches for addressing issues at specific Trigger Locations to reduce injury risk. Each row depicts an exercise across its phases of motion, highlighting the Trigger Metrics and Trigger Policies that determine when feedback is activated, alongside the corresponding feedback strategies.
This mapping underscores how expert-based policies can guide the monitoring of metrics (e.g., joint angle, motion path) at locations where injury risk is high. It also extends to personalized policies, such as monitoring muscle activity, which can vary across body types and help detect cases where the intended muscle is not engaged. Once a cue is flagged, Feedback Delivery defines the type and form of response. For example, in a lateral raise, exceeding a speed threshold of movement dynamics triggered intrinsic resistance to slow momentum, which could escalate with progressive intensity, from low to high intensity, if the deviation worsens.
By situating these elements side by side, the integrated design space highlights how different policies and feedback strategies can be tailored to the unique risks of each exercise. This connection makes explicit the logic of the system: cues sensed through body state and motion lead to feedback that constrains, guides, or withdraws in order to promote safe and effective strength training.
6. Speed Dating
To evaluate the FlexGuard design space before investing in implementation, we conducted a speed dating workshop. Following prior work using this methodology (Davidoff et al., 2007; Zhou et al., 2023; Lee et al., 2024), we presented a set of storyboards to participants with the same background from which the design space was derived. Each storyboard depicted speculative scenarios combining different feedback modalities, body locations, exercise types, sensing conditions, etc. The goal was to validate and refine the design space with the same user group from which FlexGuard was originally derived, while surfacing the strengths and limitations of different design ideas across scenarios.
We developed ten storyboard scenarios (S1-S10) that instantiated different combinations of Sensing & Triggering and Feedback Delivery dimensions (see Fig 11).
6.0.1. Study Protocol
We recruited seven beginners (B1-B7) (fewer than three months of consistent practice) to participate in three workshops (ages 24–34; 4 male, 3 female: 28M, 24M, 27F; 26M, 34M; 24M, 25F). Each workshop was conducted in our lab space, which was equipped with two pairs of dumbbells (2 and 12 pounds) and an adjustable workout bench. To establish context, participants were first asked to perform two exercises, dumbbell incline fly and lateral raise, by following short tutorial videos111https://www.youtube.com/@nutritioneering, as these movements cover two primary degrees of freedom of the shoulder. Participants were also free to try additional exercises from the tutorial set if desired. Following this exercise phase, participants were presented with a series of storyboards illustrating scenarios from our design space. They were asked to reflect on whether the scenarios aligned with their needs, expectations, and perceived challenges in strength training. Participants could continue to access the equipment and tutorial videos throughout the session. Each workshop lasted approximately one hour, with the first 10 minutes dedicated to exercise practice and the remaining 50 minutes to storyboard discussion. Participants received $20 in compensation for their time.
6.1. Results
Two researchers analyzed the workshop transcripts. The analysis surfaced three main themes that validate and extend the FlexGuard design space: (1) semantics of feedback modalities, (2) expectations for timing and progressivity of feedback, and (3) the importance of adaptability across users, exercises, and contexts. These themes highlight how participants navigated trade-offs across the storyboards. Other insights are reported together in a separate group below.
6.1.1. Semantics of Feedback Modalities
Participants emphasized that different feedback modalities carried distinct meanings in addition to their mechanical effects. intrinsic mechanisms, such as shown in S7, S8 and S9, were perceived as intuitive posture correction. As B2 noted when discussing S8, “If the material itself keeps me from going further, I immediately know what I’m doing wrong.”
extrinsiccues were instead interpreted symbolically: vibration often conveyed “something is wrong or needs attention,” prompting the user to stop and reflect. For example, in bench press (S10), if the right hand drifted ahead of the left, vibration on the right arm could signal the user to pause and allow balance to be restored before continuing. Similarly, pressure applied to a muscle group in S3 resembled a trainer’s tap to activate the correct muscle.
Inconsistent use of feedback semantics risked causing confusion, as B7 emphasized: “The feedback has to be consistent. Otherwise I won’t know what to do.” Several participants envisioned hybrid approaches that combined intrinsic mechanisms for correction with extrinsic cues for reminders or motivational prompts, highlighting the importance of designing semantics for on-body feedback.
6.1.2. Feedback Intensity Spectrum
intensityshaped how participants interpreted and responded to the feedback, with preferences spanning an entire spectrum rather than fixed levels. At the low end, participants described subtle prompts as sufficient for warnings or minor corrections where they could adjust independently. For example, in S3, participants emphasized that feedback should be gentle rather than disruptive: “It should be subtle so it doesn’t make me jump mid-movement, and once I’ve corrected it should go away” At the high end of the spectrum, participants valued strong interventions when safety or performance required it. In S5, participants appreciated the idea of feedback that actively helps complete repetitions: “I’d like this feature—it’s similar to how peers spot each other” (B5). Strong feedback was also desired for critical events such as enforcing safe ROM, as one participant noted in S8: “If I’m about to hurt myself, I want it to stop me right away.” Connecting to the concept of intrinsic feedback, B5 reflected on S7, noting how the intensity spectrum can be expressed through material properties: rigid straps provide high-intensity constraints, while elastic straps offer low-intensity guidance (Fig. 12).
6.1.3. Adaptability Across Contexts
The results on feedback semantics and intensity clearly underscore the importance of adaptability across different contexts of use. For instance, when performing an exercise, participants expressed a desire for feedback that remains subtle during movement but becomes progressively stronger as the motion deviates, favoring elastic feedback for flexibility (S7). At extrema positions, however, they preferred stronger constraints with rigid feedback to enforce safe ROM (S8). This illustrates the need for wearables that adapt not only across exercises, but also across phases of movement.
Policies such as expert-based can guide these transitions. As in the example above, knowing the thresholds for limiting ROM at both ends allows the system to escalate from elastic to rigid feedback once a defined boundary is reached. Systems must also adapt to users through personalized policies. For example, when fatigue is detected, the system can provide supportive assistance (S5). In S3, if the exercise changes and the target region shifts, the system should adapt reminders to activate the appropriate muscle group. Adaptability also extends to user expertise: as one participant noted, “Once I’ve learned and memorized the correct movement, a reminder might be less necessary.”
6.1.4. Other Findings
Some scenarios, such as S2, were seen as potentially useful in specific contexts. For example, when training with a mild injury, stabilization could help bear the load from the affected muscles.
At the same time, participants reflected on how wearables might influence learning. Several worried that excessive support could reduce opportunities to internalize correct form. For instance, participants raised concerns about S2: on one hand, users should learn through practice, but on the other, they appreciated how the wearable could provide stabilization to ensure the correct muscles were engaged. In contrast, S9 was regarded as more supportive of learning by providing feedback on tempo. As B1 noted, “I could know if I go too fast or not, because I think that’s important to know.”
Participants also expressed concerns about potential risks with locking mechanisms such as in S6. If weights were falling and users were unable to reach a safe position while locked, this could be dangerous. As B5 explained, “I will hold myself accountable and not rush through the steps, so I might not use it.” To address this, B3 suggested an “emergency button” that would release the lock when things go wrong.
6.1.5. Summary
Together, these results validate the design space dimensions presented in FlexGuard while also surfacing opportunities for refinement. Because the speed dating method relied on storyboards, participants’ reflections were necessarily speculative rather than grounded in embodied practice despite trying a few exercises in the lab. Nevertheless, adaptability emerged as a recurring priority, particularly for feedback that adjusts across phases of movement. The workshops also clarified the scope of what should be included within the design space. For example, we were initially uncertain whether feedback was necessary for inter-limb coordination (S10), since users can naturally perceive weight imbalance during an exercise. However, across all sessions, participants (B1, B2, B5, B7) consistently emphasized the value of on-body cues to confirm user’s interpretation of the situation—B5: “Although I can feel something is off, the vibration tells me exactly what’s happening.”
7. Workshops with a Proof-of-Concept
While extrinsic modalities have been more extensively studied in training contexts (Spelmezan et al., 2009; Jo et al., 2023b; Huang et al., 2025; Elvitigala et al., 2019), intrinsic feedback remains comparatively underexplored and more difficult to adapt dynamically. We therefore focused our follow-up study on investigating the adaptability of intrinsic feedback in practice, providing finer-grained validation of how design space principles translate into embodied exercise contexts.
Building on the findings of the speed dating study, we conducted a workshop with a proof-of-concept (PoC) to investigate how adaptability in intrinsic feedback could be experienced in practice.
7.1. PoC Design
The PoC system approximated a closed loop of sensing, decision-making, and actuation. Two external cameras captured front and side views of participants’ movements, which were processed on a server for real-time pose estimation and compared against ground-truth postures derived from tutorial videos. The system flagged form deviations and signaled adjustments to strap intensity (Fig. 13).
The digital component of the PoC (Fig. 14) consisted of a user interface and a pose-estimation backend. The interface presented participants with a demonstration of the correct form for each exercise, while the backend was powered by MediaPipe joint tracking for real-time estimation of joint positions and angles. Using both front- and side-view cameras, the system primarily tracked key joints relevant to the exercises, and calculated their current angles relative to the target posture.
The PoC suit prototype consisted of a skin-fitting shirt equipped with rigid straps across the chest, shoulders, upper arms, and posterior forearms, complemented by elastic straps on the shoulders. The PoC suit was designed around lateral raise and incline dumbbell fly (Fig. 15), as these exercises together span two key degrees of freedom of the shoulder, one of the most common sites of injury in strength training (Hedrick and Wada, 2008). To account for phase adaptation, each strap was mapped to different phases of movement. Although strap adjustments were performed manually by researchers, this setup simulated how a future motorized system could dynamically modulate intrinsic resistance in real time.
For the incline dumbbell fly, r1 supported the execution phase by guiding the elbow from a folded to a straightened position, and r3 reinforced the completion phase posture by maintaining upright torso alignment. Strap r4 provided feedback on the folding limit of the elbow, preventing unsafe flexion and indicating the appropriate end-point of elbow bend. In parallel, l1 offered progressive resistance throughout the motion, ensuring a smooth and safe trajectory of the arms. In the lateral raise, r2 regulated the arm elevation phase by linking torso and shoulder movement, while r4 again controlled elbow flexion, restricting collapse and guiding the fold to its safe range.
7.2. Protocol
We recruited six beginners in strength training (fewer than three months of consistent training; 24M, 23M, 29M, 21F, 24F, 24M). The study was conducted in our lab, which was equipped with an adjustable workout bench and two sets of dumbbells. After providing informed consent, participants put on the PoC suit, with disposable elbow sleeves supplied to prevent direct skin contact. At least one researcher of the same gender as the participant assisted with putting on the suit and adjusting the straps. The session proceeded as follows:
Calibration: For each participant, we established baseline “low” and “high” feedback intensity levels and recorded their locations on the straps. We also measured rigid strap tension using a Newton spring scale.
Exercise Trials: Participants performed five repetitions of each exercise while following a tutorial video displayed on the monitor of the PoC platform. The PoC tracked the exercise phase-by-phase. When posture deviations occurred, researchers manually adjusted the straps to simulate adaptive feedback.
Repeated Trials: Participants repeated the exercise for another five repetitions at the adjusted intensity level to experience the adjusted feedback intensity.
Survey and Interview: After the PoC trial, participants completed a questionnaire on perceived support, comfort, and appropriateness of feedback. While completing the questionnaire, they were asked to think aloud, and we conducted follow-up questions to further explore their reflections on adaptability, effectiveness, and potential improvements.
Each study session lasted approximately one hour, and participants received $20 in compensation for their time and effort.
7.3. Results
We analyzed the survey responses, and two researchers reviewed the interview recordings, grouping the findings into themes. Our results further validated the FlexGuard design space, particularly its emphasis on phase-specific movement. Figure 16 summarizes participants’ ratings across survey items (1–7 Likert scale), showing generally positive perceptions of on-body feedback aligned with different phases of movement.
7.3.1. Phase-Specific Adaptations
Participants strongly valued feedback that adapted to different phases of movement. Survey responses reflected this, with high ratings for “Feedback was appropriately aligned with different phases of movement” (mean = 5.7). As P1 confirmed, “I can properly feel it in the different phases as I move.” Similarly, P3 emphasized the importance of differential support: “When lowering the dumbbells, I need more resistance, but at the top I don’t want to feel held back.” Together, these accounts reinforce phase-specific adaptability as both perceptible and appreciated, establishing it as a core design principle.
7.3.2. Correctness vs. Flexibility
A recurring theme was the balance between enforcing correctness and allowing flexibility. The PoC was rated highly for corrective support (mean = 6.0) but lower for comfort (mean = 4.2). Higher feedback intensity offered stronger prevention against incorrect posture but came at the cost of reduced comfort, as reflected in the lower survey ratings for intensity appropriateness, with some participants noting that it felt too strong. Participants appreciated the rigid feedback that restricted unsafe ROM, but also valued elastic support that permitted controlled deviation. As P5 noted, “If it’s too strict, I feel like I can’t really move.”
However, participants also reported that the feedback from the elastic straps was sometimes too subtle, especially when near the intended plane of motion. This limitation contributed to a lower score on the item about staying on the correct movement path (mean = 4.3), as the elastic straps were not calibrated due to equipment constraints. Nevertheless, participants still felt guided during exercise (mean = 5.3), noting that the elastic straps provided a “kickback” force that pushed their arms back on track when they drifted too high or too low from the recommended plane.
7.3.3. Supporting Injury Prevention
Participants recognized the system’s potential to support injury prevention. Participants rated highly on “The suit helped me avoid unsafe ranges of motion” (mean = 5.7). P2 explained, “This suit stops me before going too far.” Similarly, P4 emphasized the value of strong corrective cues: “The suit stops me right away, which is really great if I’m about to hurt myself.” These results validate the role of rigid straps as hard stops for critical safety.
7.3.4. Suggestions and Improvements
While overall perceptions were positive, participants also identified areas for improvement. Comfort was a recurring concern, with several noting issues such as ill-fitting shirt sizes and strap tightness that could limit long-term use. As P3 remarked, “I like the idea, but I wouldn’t want to wear this for a whole workout—it gets uncomfortable. Maybe because the shirt is size M.” P6 also observed that when strap intensity increased, tension was transferred to anchoring points, causing discomfort in the forearm even on sides not directly involved. These reflections suggest opportunities for refining future prototypes to improve fit, comfort, and localized feedback delivery.
8. Discussion
8.1. Implications for Sports and Fitness Wearables
By situating injury prevention within a design space framework, FlexGuard advances the conversation on how future sports and fitness wearables can evolve from passive tracking systems to interactive training partners. Our findings illustrate how wearables can move beyond performance monitoring to actively guide movement, prevent injury, and support motor learning. This reframing has broader implications for adjacent domains such as rehabilitation, physical education, and recreational training, where embodied safety and skill acquisition are equally central. In this way, FlexGuard demonstrates how design spaces can serve as actionable frameworks for envisioning the next generation of intelligent, interactive wearables.
8.2. Intelligence in Materials
Our findings show that there is a need for greater intelligence in materials themselves. Many participant needs pointed toward adaptive intrinsic feedback, where feedback is conveyed through material properties such as rigidity or elasticity. Unlike extrinsic actuation, which can be modulated digitally or mechanically, intrinsic feedback relies on the physical response of the material. To make such feedback adaptive, materials themselves must become “smart” and dynamically tunable. This suggests a path forward through mechano- or physical intelligence (Wang and Li, 2023), where material properties are engineered to support highly adaptive, wearable use cases.
8.3. Balancing Corrective and Supportive Feedback
There is a tension between feedback that is corrective and feedback that is supportive. Participants frequently contrasted rigid straps, which acted as “hard stops,” with elastic straps, which functioned more as guides. This mirrors pedagogical debates in coaching, where overcorrection may undermine motor learning while undercorrection risks unsafe execution. The implication for wearable design is the need for hybrid strategies that balance these roles, such as graduated resistance that escalates when errors persist or supportive cues that reinforce correct posture without being overbearing. Future sports and fitness wearables may need to embody this balance to ensure both safety and long-term skill acquisition.
9. Limitations and Future Work
Our work carries several limitations that inform directions for future research. First, we investigated only upper-body movements. While these capture many common injury-prone patterns, they represent only a subset of the body. Some insights may generalize to other regions with similar roles, such as joints that govern ROM, supporting muscle groups, and related structures, but future work should extend our approach to lower-body and full-body exercises, where different biomechanical demands may surface new requirements for sensing and feedback design.
Second, in our PoC validation we relied on manual configuration of straps to set feedback intensity levels. While sufficient for this study, this approach introduced setup delays and required facilitator involvement. Although strength training movements are relatively straightforward, minimizing delays will be especially important when studying on-body feedback in complex, time-sensitive movements. It also remains underexplored how multiple feedback cues delivered in a short time window can be coordinated to minimize cognitive load.
Third, our strap tension measurements (7–15 N across different locations) revealed an important limitation for wearable systems relying on passive, strap-based feedback. Stronger corrective cues require higher strap tensions, yet higher tensions make it more difficult to maintain comfort and stability without bulky motors or anchoring structures. Designing garments that can distribute forces evenly across the body while remaining flexible and wearable in dynamic exercise contexts remains an open challenge.
Future implementations should build on this foundation by automating configuration through adaptive calibration, onboard sensing, or machine learning to reduce setup burden and enable seamless, on-the-fly personalization. Such systems should remain lightweight and body-conforming to support strength training. This direction would move beyond design-space validation toward deployable wearable systems capable of delivering continuous, real-time guidance in everyday training settings.
10. Conclusion
This paper presented FlexGuard, a design space for on-body wearable feedback for injury prevention in strength training derived from nine co-design workshops with trainer–trainee pairs using a technology probe that explored both extrinsic and intrinsic modalities. FlexGuard is articulated through dimensions of Sensing & Triggering and Feedback Delivery, providing designers with a framework to associate injury risks in specific muscles with appropriate forms of feedback to prevent injuries during exercise. We validated this design space through a speed dating study with seven beginners, which highlighted the importance of designing semantics when using extrinsic modalities, the timing of feedback delivery, and adaptation across fast-changing training contexts. We further validated FlexGuard at the movement level through a PoC prototype evaluated with six beginners, where results revealed tensions between effectiveness and comfort, variability in perception across users, and the challenges of achieving dynamically adaptive feedback. Overall, our findings show that FlexGuard is a valid and useful framework for designing on-body feedback systems aimed at injury prevention in strength training.
Acknowledgements.
We thank all of the trainers and trainees who participated in our study and shared their time and expertise.References
- (1)
- Al Maimani and Roudaut (2017) Ahmed Al Maimani and Anne Roudaut. 2017. Frozen Suit: Designing a Changeable Stiffness Suit and its Application to Haptic Games. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 2440–2448. doi:10.1145/3025453.3025655
- Aliyana et al. (2024) Akshaya Kumar Aliyana, Danying Yang, Orathai Tangsirinaruenart, and George K. Stylios. 2024. A garment-integrated textile stitch-based strain sensor device, IoT-Enabled for enhanced wearable sportswear applications. Results in Engineering 23 (2024), 102794. doi:10.1016/j.rineng.2024.102794
- Alqarni (2019) Ahmed Mohammed Alqarni. 2019. Common injuries in resistance training. Saudi Journal of Sports Medicine 19, 2 (2019), 38–42.
- Alu et al. (2025) Andrea Alu, Andres F Arrieta, Emanuela Del Dottore, Michael D Dickey, Samuele Ferracin, Ryan L Harne, Helmut Hauser, Qiguang He, Jonathan Brigham Hopkins, Lance Hyatt, Suyi Li, Stefano Mariani, Barbara Mazzolai, Alessio Mondini, Aniket Pal, Daniel J Preston, Anoop Rajappan, Jordan R. Raney, Pedro Reis, Stephen Andrew Sarles, Metin Sitti, Uba K. Ubamanyu, Martin van Hecke, and Kon-Well Wang. 2025. Roadmap on embodying mechano-intelligence and computing in functional materials and structures. Smart Materials and Structures (2025). http://iopscience.iop.org/article/10.1088/1361-665X/adb7aa
- Amershi et al. (2014) Saleema Amershi, Maya Cakmak, W. Bradley Knox, and Todd Kulesza. 2014. Power to the People: The Role of Humans in Interactive Machine Learning. AI Magazine 35, 4 (2014), 105–120. doi:10.1609/aimag.v35i4.2513 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1609/aimag.v35i4.2513
- Aslam et al. (2025) Sumaira Aslam, Jean De Dieu Habyarimana, and Shi Yong Bin. 2025. Neuromuscular adaptations to resistance training in elite versus recreational athletes. Frontiers in Physiology 16 (2025), 1598149.
- Barnes et al. (2021) Matthew J Barnes, Ashley Petterson, and Darryl J Cochrane. 2021. Peak power output and onset of muscle activation during high pull exercise. The Journal of Strength & Conditioning Research 35, 3 (2021), 675–679.
- Braun and Clarke (2019) Virginia Braun and Victoria Clarke. 2019. Reflecting on reflexive thematic analysis. Qualitative Research in Sport, Exercise and Health 11, 4 (2019), 589–597. doi:10.1080/2159676X.2019.1628806 arXiv:https://doi.org/10.1080/2159676X.2019.1628806
- Brooks et al. (2024) Jas Brooks, Alex Mazursky, Janice Hixon, and Pedro Lopes. 2024. Augmented Breathing via Thermal Feedback in the Nose. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 26, 11 pages. doi:10.1145/3654777.3676438
- Caserman et al. (2021) Polona Caserman, Clemens Krug, and Stefan Göbel. 2021. Recognizing full-body exercise execution errors using the teslasuit. Sensors 21, 24 (2021), 8389.
- Catapult (2025) Catapult. 2025. Capture, Analyze, Plan, Share every aspect of performance. Retrieved Jul 15, 2025 from https://www.catapult.com
- Chang et al. (2020) Zekun Chang, Tung D. Ta, Koya Narumi, Heeju Kim, Fuminori Okuya, Dongchi Li, Kunihiro Kato, Jie Qi, Yoshinobu Miyamoto, Kazuya Saito, and Yoshihiro Kawahara. 2020. Kirigami Haptic Swatches: Design Methods for Cut-and-Fold Haptic Feedback Mechanisms. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3313831.3376655
- Davidoff et al. (2007) Scott Davidoff, Min Kyung Lee, Anind K Dey, and John Zimmerman. 2007. Rapidly exploring application design through speed dating. In International conference on ubiquitous computing. Springer, 429–446.
- Delazio et al. (2018) Alexandra Delazio, Ken Nakagaki, Roberta L. Klatzky, Scott E. Hudson, Jill Fain Lehman, and Alanson P. Sample. 2018. Force Jacket: Pneumatically-Actuated Jacket for Embodied Haptic Experiences. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (Montreal QC, Canada) (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3173574.3173894
- Elvitigala et al. (2024) Don Samitha Elvitigala, Armağan Karahanoğlu, Andrii Matviienko, Laia Turmo Vidal, Dees Postma, Michael D Jones, Maria F. Montoya, Daniel Harrison, Lars Elbæk, Florian Daiber, Lisa Anneke Burr, Rakesh Patibanda, Paolo Buono, Perttu Hämäläinen, Robby Van Delden, Regina Bernhaupt, Xipei Ren, Vincent Van Rheden, Fabio Zambetta, Elise Van Den Hoven, Carine Lallemand, Dennis Reidsma, and Florian ‘Floyd’ Mueller. 2024. Grand Challenges in SportsHCI. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 312, 20 pages. doi:10.1145/3613904.3642050
- Elvitigala et al. (2019) Don Samitha Elvitigala, Denys J.C. Matthies, Löic David, Chamod Weerasinghe, and Suranga Nanayakkara. 2019. GymSoles: Improving Squats and Dead-Lifts by Visualizing the User’s Center of Pressure. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3290605.3300404
- Escudero (2024) Julio Andres Escudero. 2024. Haptic Scores: Non-Visual Ways of Experiencing Choreography. In Companion of the 2024 on ACM International Joint Conference on Pervasive and Ubiquitous Computing (Melbourne VIC, Australia) (UbiComp ’24). Association for Computing Machinery, New York, NY, USA, 281–285. doi:10.1145/3675094.3678368
- Esposito et al. (2018) Daniele Esposito, Emilio Andreozzi, Antonio Fratini, Gaetano D Gargiulo, Sergio Savino, Vincenzo Niola, and Paolo Bifulco. 2018. A Piezoresistive Sensor to Measure Muscle Contraction and Mechanomyography. Sensors 18, 8 (2018). doi:10.3390/s18082553
- Evke et al. (2019) Erin E Evke, Dilara Meli, and Max Shtein. 2019. Developable rotationally symmetric Kirigami-based structures as sensor platforms. Advanced Materials Technologies 4, 12 (2019), 1900563.
- Faigenbaum and Myer (2010) Avery D Faigenbaum and Gregory D Myer. 2010. Resistance training among young athletes: safety, efficacy and injury prevention effects. British journal of sports medicine 44, 1 (2010), 56–63.
- Fisher et al. (2023) James Fisher, Patroklos Androulakis-Korakakis, Jürgen Giessing, Eric Helms, Brad Schoenfeld, Dave Smith, and Richard Winett. 2023. Supervision during resistance training: A comparison of trainer and trainee perceptions. International Journal of Strength and Conditioning 3, 1 (2023).
- Fisher (2025) James P Fisher. 2025. Supervision During Strength Training—the Interplay with Facilitation, Feedback and Attentional Focus: A Narrative Review. Sports Medicine (2025), 1–9.
- Gabbett (2020) Tim J Gabbett. 2020. How much? How fast? How soon? Three simple concepts for progressing training loads to minimize injury risk and enhance performance. Journal of orthopaedic & sports physical therapy 50, 10 (2020), 570–573.
- Gillies (2019) Marco Gillies. 2019. Understanding the Role of Interactive Machine Learning in Movement Interaction Design. ACM Trans. Comput.-Hum. Interact. 26, 1, Article 5 (Feb. 2019), 34 pages. doi:10.1145/3287307
- Günther et al. (2019) Sebastian Günther, Mohit Makhija, Florian Müller, Dominik Schön, Max Mühlhäuser, and Markus Funk. 2019. PneumAct: Pneumatic Kinesthetic Actuation of Body Joints in Virtual Reality Environments. In Proceedings of the 2019 on Designing Interactive Systems Conference (San Diego, CA, USA) (DIS ’19). Association for Computing Machinery, New York, NY, USA, 227–240. doi:10.1145/3322276.3322302
- Hamdan et al. (2019) Nur Al-huda Hamdan, Adrian Wagner, Simon Voelker, Jürgen Steimle, and Jan Borchers. 2019. Springlets: Expressive, Flexible and Silent On-Skin Tactile Interfaces. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–14. doi:10.1145/3290605.3300718
- Han et al. (2023) Violet Yinuo Han, Abena Boadi-Agyemang, Yuyu Lin, David Lindlbauer, and Alexandra Ion. 2023. Parametric Haptics: Versatile Geometry-based Tactile Feedback Devices. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (San Francisco, CA, USA) (UIST ’23). Association for Computing Machinery, New York, NY, USA, Article 65, 13 pages. doi:10.1145/3586183.3606766
- Hartmann et al. (2007) Björn Hartmann, Leith Abdulla, Manas Mittal, and Scott R. Klemmer. 2007. Authoring sensor-based interactions by demonstration with direct manipulation and pattern recognition. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’07). Association for Computing Machinery, New York, NY, USA, 145–154. doi:10.1145/1240624.1240646
- Hedrick and Wada (2008) Allen Hedrick and Hiroaki Wada. 2008. Weightlifting movements: do the benefits outweigh the risks? Strength & Conditioning Journal 30, 6 (2008), 26–35.
- Herbaut and Tuloup (2025) Alexis Herbaut and Edouard Tuloup. 2025. Effect of weightlifting belts on lumbar biomechanics and muscle activity in deadlift and squat. Sports Engineering 28, 1 (2025), 1–9.
- Hu et al. (2025) Hongci Hu, Mengqi Jiang, Kai Lin, Kinor Shou-xiang Jiang, and Ziqian Bai. 2025. ReKnit-Care: A Seamless-Knitted Sensing Glove for Sensory Rehabilitation and Adaptive Haptic Feedback. In Proceedings of the Nineteenth International Conference on Tangible, Embedded, and Embodied Interaction (TEI ’25). Association for Computing Machinery, New York, NY, USA, Article 58, 7 pages. doi:10.1145/3689050.3705979
- Huang et al. (2025) Bingjian Huang, Siyi Ren, Yuewen Luo, Qilong Cheng, Hanfeng Cai, Yeqi Sang, Mauricio Sousa, Paul H Dietz, and Daniel Wigdor. 2025. VibraForge: A Scalable Prototyping Toolkit For Creating Spatialized Vibrotactile Feedback Systems. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 1137, 18 pages. doi:10.1145/3706598.3714273
- Hutchinson et al. (2003) Hilary Hutchinson, Wendy Mackay, Bo Westerlund, Benjamin B. Bederson, Allison Druin, Catherine Plaisant, Michel Beaudouin-Lafon, Stéphane Conversy, Helen Evans, Heiko Hansen, Nicolas Roussel, and Björn Eiderbäck. 2003. Technology probes: inspiring design for and with families. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Ft. Lauderdale, Florida, USA) (CHI ’03). Association for Computing Machinery, New York, NY, USA, 17–24. doi:10.1145/642611.642616
- Islam et al. (2022) Md Shafiqul Islam, Sang Won Lee, Samantha M. Harden, and Sol Lim. 2022. Effects of vibrotactile feedback on yoga practice. Frontiers in Sports and Active Living Volume 4 - 2022 (2022). doi:10.3389/fspor.2022.1005003
- Jo et al. (2023a) Hye-Young Jo, Laurenz Seidel, Michel Pahud, Mike Sinclair, and Andrea Bianchi. 2023a. FlowAR: How Different Augmented Reality Visualizations of Online Fitness Videos Support Flow for At-Home Yoga Exercises. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 469, 17 pages. doi:10.1145/3544548.3580897
- Jo et al. (2023b) Hye-Young Jo, Chan Hu Wie, Yejin Jang, Dong-Uk Kim, Yurim Son, and Yoonji Kim. 2023b. TrainerTap: Weightlifting Support System Prototype Simulating Personal Trainer’s Tactile and Auditory Guidance. In Adjunct Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (San Francisco, CA, USA) (UIST ’23 Adjunct). Association for Computing Machinery, New York, NY, USA, Article 14, 3 pages. doi:10.1145/3586182.3616644
- Johnson et al. (2023) Tate Johnson, Dinesh K Patel, Humphrey Yang, Umut Serdar Civici, Adriane Fernandes Minori, and Lining Yao. 2023. FlexTure: Designing Configurable and Dynamic Surface Features. In Proceedings of the 2023 ACM Designing Interactive Systems Conference (Pittsburgh, PA, USA) (DIS ’23). Association for Computing Machinery, New York, NY, USA, 580–593. doi:10.1145/3563657.3595995
- Kao et al. (2018) Hsin-Liu (Cindy) Kao, Miren Bamforth, David Kim, and Chris Schmandt. 2018. Skinmorph: texture-tunable on-skin interface through thin, programmable gel. In Proceedings of the 2018 ACM International Symposium on Wearable Computers (Singapore, Singapore) (ISWC ’18). Association for Computing Machinery, New York, NY, USA, 196–203. doi:10.1145/3267242.3267262
- Kaul and Rohs (2017) Oliver Beren Kaul and Michael Rohs. 2017. HapticHead: A Spherical Vibrotactile Grid around the Head for 3D Guidance in Virtual and Augmented Reality. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 3729–3740. doi:10.1145/3025453.3025684
- Keelawat and Suzuki (2024) Panayu Keelawat and Ryo Suzuki. 2024. Transforming Procedural Instructions into In-Situ Augmented Reality Guides with InstructAR. In Adjunct Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST Adjunct ’24). Association for Computing Machinery, New York, NY, USA, Article 70, 3 pages. doi:10.1145/3672539.3686321
- Kilic Afsar et al. (2021) Ozgun Kilic Afsar, Ali Shtarbanov, Hila Mor, Ken Nakagaki, Jack Forman, Karen Modrei, Seung Hee Jeong, Klas Hjort, Kristina Höök, and Hiroshi Ishii. 2021. OmniFiber: Integrated Fluidic Fiber Actuators for Weaving Movement based Interactions into the ‘Fabric of Everyday Life’. In The 34th Annual ACM Symposium on User Interface Software and Technology (Virtual Event, USA) (UIST ’21). Association for Computing Machinery, New York, NY, USA, 1010–1026. doi:10.1145/3472749.3474802
- Kim and Asbeck (2022) Hubert Kim and Alan T. Asbeck. 2022. The Effects of Torque Magnitude and Stiffness in Arm Guidance Through Joint Torque Feedback. IEEE Access 10 (2022), 5842–5854. doi:10.1109/ACCESS.2022.3141981
- Knoblauch et al. (2014) Hubert Knoblauch, René T., and Bernt Schnettler. 2014. Video Analysis and Videography Qualitative Methods.
- Kosmalla et al. (2016) Felix Kosmalla, Frederik Wiehr, Florian Daiber, Antonio Krüger, and Markus Löchtefeld. 2016. ClimbAware: Investigating Perception and Acceptance of Wearables in Rock Climbing. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’16). Association for Computing Machinery, New York, NY, USA, 1097–1108. doi:10.1145/2858036.2858562
- Kowsar et al. (2019) Yousef Kowsar, Eduardo Velloso, Lars Kulik, and Christopher Leckie. 2019. LiftSmart: a monitoring and warning wearable for weight trainers. In Adjunct Proceedings of the 2019 ACM International Joint Conference on Pervasive and Ubiquitous Computing and Proceedings of the 2019 ACM International Symposium on Wearable Computers (London, United Kingdom) (UbiComp/ISWC ’19 Adjunct). Association for Computing Machinery, New York, NY, USA, 298–301. doi:10.1145/3341162.3343795
- Kwon et al. (2022) Junghan Kwon, Inrak Choi, Myungsun Park, Jeongin Moon, Bomin Jeong, Prabhat Pathak, Jooeun Ahn, and Yong-Lae Park. 2022. Selectively stiffening garments enabled by cellular composites. Advanced Materials Technologies 7, 9 (2022), 2101543.
- Latella et al. (2024) Christopher Latella, Joel Garrett, and Daniel van den Hoek. 2024. How technological impacts on performance have been managed in elite sport: a powerlifting example. Journal of Applied Physiology 137, 4 (2024), 821–822.
- Latella et al. (2019) Christopher Latella, Daniel Van Den Hoek, and Wei-Peng Teo. 2019. Differences in strength performance between novice and elite athletes: Evidence from powerlifters. The Journal of Strength & Conditioning Research 33 (2019), S103–S112.
- Lee et al. (2024) Soohwan Lee, Seoyeong Hwang, Ian Oakley, and Kyungho Lee. 2024. Expanding the Design Space of Vision-based Interactive Systems for Group Dance Practice. In Proceedings of the 2024 ACM Designing Interactive Systems Conference (Copenhagen, Denmark) (DIS ’24). Association for Computing Machinery, New York, NY, USA, 2768–2787. doi:10.1145/3643834.3661568
- Leiva et al. (2021) Germán Leiva, Jens Emil Grønbæk, Clemens Nylandsted Klokmose, Cuong Nguyen, Rubaiat Habib Kazi, and Paul Asente. 2021. Rapido: Prototyping Interactive AR Experiences through Programming by Demonstration. In The 34th Annual ACM Symposium on User Interface Software and Technology (Virtual Event, USA) (UIST ’21). Association for Computing Machinery, New York, NY, USA, 626–637. doi:10.1145/3472749.3474774
- Li et al. (2017) Shuguang Li, Daniel M. Vogt, Daniela Rus, and Robert J. Wood. 2017. Fluid-driven origami-inspired artificial muscles. Proceedings of the National Academy of Sciences 114, 50 (2017), 13132–13137. doi:10.1073/pnas.1713450114 arXiv:https://www.pnas.org/doi/pdf/10.1073/pnas.1713450114
- Li and Luo (2024) Yunzhu Li and Yiyue Luo. 2024. Intelligent textiles are looking bright. Science 384, 6691 (2024), 29–30. doi:10.1126/science.ado5922 arXiv:https://www.science.org/doi/pdf/10.1126/science.ado5922
- Lin et al. (2024a) Yuyu Lin, Jesse T Gonzalez, Zhitong Cui, Yash Rajeev Banka, and Alexandra Ion. 2024a. ConeAct: A Multistable Actuator for Dynamic Materials. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 324, 16 pages. doi:10.1145/3613904.3642949
- Lin et al. (2025a) Yuyu Lin, Hatice Gokcen Guner, Jianzhe Gu, Sonia Prashant, and Alexandra Ion. 2025a. Wearable Material Properties: Passive Wearable Microstructures as Adaptable Interfaces for the Physical Environment. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 305, 16 pages. doi:10.1145/3706598.3714215
- Lin et al. (2024b) Yilong Lin, Peng Zhang, Eyal Ofek, and Seungwoo Je. 2024b. ArmDeformation: Inducing the Sensation of Arm Deformation in Virtual Reality Using Skin-Stretching. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 407, 18 pages. doi:10.1145/3613904.3642518
- Lin et al. (2025b) Yuyu Lin, Dian Zhu, Anoushka Naidu, Kenneth Yu, Deon Harper, Eni Halilaj, Douglas Weber, Deborah Ellen Kenney, Adam J. Popchak, Mark Baratz, and Alexandra Ion. 2025b. Personalized Bistable Orthoses for Rehabilitation of Finger Joints. In Proceedings of the 38th Annual ACM Symposium on User Interface Software and Technology (Busan, Republic of Korea) (UIST ’25). Association for Computing Machinery, New York, NY, USA. doi:10.1145/3746059.3747643
- Lu et al. (2019) Qiuyu Lu, Jifei Ou, João Wilbert, André Haben, Haipeng Mi, and Hiroshi Ishii. 2019. milliMorph – Fluid-Driven Thin Film Shape-Change Materials for Interaction Design. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (New Orleans, LA, USA) (UIST ’19). Association for Computing Machinery, New York, NY, USA, 663–672. doi:10.1145/3332165.3347956
- Luo et al. (2024) Yiyue Luo, Chao Liu, Young Joong Lee, Joseph DelPreto, Kui Wu, Michael Foshey, Daniela Rus, Tomás Palacios, Yunzhu Li, Antonio Torralba, et al. 2024. Adaptive tactile interaction transfer via digitally embroidered smart gloves. Nature communications 15, 1 (2024), 868.
- Ma et al. (2024) Dizhi Ma, Xiyun Hu, Jingyu Shi, Mayank Patel, Rahul Jain, Ziyi Liu, Zhengzhe Zhu, and Karthik Ramani. 2024. avaTTAR: Table Tennis Stroke Training with Embodied and Detached Visualization in Augmented Reality. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 35, 16 pages. doi:10.1145/3654777.3676400
- Machek et al. (2021) Steven B Machek, Thomas D Cardaci, Dylan T Wilburn, Mitchell C Cholewinski, Scarlett Lin Latt, Dillon R Harris, and Darryn S Willoughby. 2021. Neoprene knee sleeves of varying tightness augment barbell squat one repetition maximum performance without improving other indices of muscular strength, power, or endurance. The Journal of Strength & Conditioning Research 35 (2021), S6–S15.
- Mahadevan et al. (2022) Karthik Mahadevan, Yan Chen, Maya Cakmak, Anthony Tang, and Tovi Grossman. 2022. Mimic: In-Situ Recording and Re-Use of Demonstrations to Support Robot Teleoperation. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology (Bend, OR, USA) (UIST ’22). Association for Computing Machinery, New York, NY, USA, Article 40, 13 pages. doi:10.1145/3526113.3545639
- Mahadevan et al. (2024) Karthik Mahadevan, Jonathan Chien, Noah Brown, Zhuo Xu, Carolina Parada, Fei Xia, Andy Zeng, Leila Takayama, and Dorsa Sadigh. 2024. Generative Expressive Robot Behaviors using Large Language Models. In Proceedings of the 2024 ACM/IEEE International Conference on Human-Robot Interaction (Boulder, CO, USA) (HRI ’24). Association for Computing Machinery, New York, NY, USA, 482–491. doi:10.1145/3610977.3634999
- Morris et al. (2014) Dan Morris, T. Scott Saponas, Andrew Guillory, and Ilya Kelner. 2014. RecoFit: using a wearable sensor to find, recognize, and count repetitive exercises. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Toronto, Ontario, Canada) (CHI ’14). Association for Computing Machinery, New York, NY, USA, 3225–3234. doi:10.1145/2556288.2557116
- Paay et al. (2020) Jeni Paay, Jesper Kjeldskov, Frederick Sorensen, Thomas Jensen, and Oren Tirosh. 2020. Weight-Mate: Adaptive Training Support for Weight Lifting. In Proceedings of the 31st Australian Conference on Human-Computer-Interaction (Fremantle, WA, Australia) (OzCHI ’19). Association for Computing Machinery, New York, NY, USA, 95–105. doi:10.1145/3369457.3369466
- Pardomuan et al. (2024) Jefferson Pardomuan, Shio Miyafuji, Nobuhiro Takahashi, and Hideki Koike. 2024. VabricBeads: Variable Stiffness Structured Fabric using Artificial Muscle in Woven Beads. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 335, 17 pages. doi:10.1145/3613904.3642401
- Park et al. (2020) Junghoon Park, Junhwan Choi, Sangjoon J. Kim, Kap-Ho Seo, and Jung Kim. 2020. Design of an Inflatable Wrinkle Actuator With Fast Inflation/Deflation Responses for Wearable Suits. IEEE Robotics and Automation Letters 5, 3 (2020), 3799–3805. doi:10.1109/LRA.2020.2976299
- Payne et al. (2021) William Christopher Payne, Yoav Bergner, Mary Etta West, Carlie Charp, R. Benjamin Shapiro, Danielle Albers Szafir, Edd V. Taylor, and Kayla DesPortes. 2021. danceON: Culturally Responsive Creative Computing. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 96, 16 pages. doi:10.1145/3411764.3445149
- Peiris et al. (2017) Roshan Lalintha Peiris, Wei Peng, Zikun Chen, Liwei Chan, and Kouta Minamizawa. 2017. ThermoVR: Exploring Integrated Thermal Haptic Feedback with Head Mounted Displays. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 5452–5456. doi:10.1145/3025453.3025824
- Pu et al. (2025) Kevin Pu, Daniel Lazaro, Ian Arawjo, Haijun Xia, Ziang Xiao, Tovi Grossman, and Yan Chen. 2025. Assistance or Disruption? Exploring and Evaluating the Design and Trade-offs of Proactive AI Programming Support. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 152, 21 pages. doi:10.1145/3706598.3713357
- Rao et al. (2024) SK Rao, R Savalgi, VK Suresh, N Pyati, and S Krishnamurthy. 2024. FormFit: A physical exercise form correction system using BlazePose. In Computer Science Engineering. CRC Press, 490–497.
- Rudolph et al. (2022) Julius Cosmo Romeo Rudolph, David Holman, Bruno De Araujo, Ricardo Jota, Daniel Wigdor, and Valkyrie Savage. 2022. Sensing Hand Interactions with Everyday Objects by Profiling Wrist Topography. In Proceedings of the Sixteenth International Conference on Tangible, Embedded, and Embodied Interaction (Daejeon, Republic of Korea) (TEI ’22). Association for Computing Machinery, New York, NY, USA, Article 14, 14 pages. doi:10.1145/3490149.3501320
- Saini et al. (2024) Aryan Saini, Rakesh Patibanda, Nathalie Overdevest, Elise Van Den Hoven, and Florian ‘Floyd’ Mueller. 2024. PneuMa: Designing Pneumatic Bodily Extensions for Supporting Movement in Everyday Life. In Proceedings of the Eighteenth International Conference on Tangible, Embedded, and Embodied Interaction (Cork, Ireland) (TEI ’24). Association for Computing Machinery, New York, NY, USA, Article 1, 16 pages. doi:10.1145/3623509.3633349
- Seçkin et al. (2023) Ahmet Çağdaş Seçkin, Bahar Ateş, and Mine Seçkin. 2023. Review on Wearable Technology in sports: Concepts, Challenges and opportunities. Applied sciences 13, 18 (2023), 10399.
- Semeraro and Turmo Vidal (2022) Alessandra Semeraro and Laia Turmo Vidal. 2022. Visualizing Instructions for Physical Training: Exploring Visual Cues to Support Movement Learning from Instructional Videos. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 90, 16 pages. doi:10.1145/3491102.3517735
- Snouzy (2025) Snouzy. 2025. Modern open‐source fitness coaching platform. Create workout plans, track progress, and access a comprehensive exercise database. https://www.workout.cool/. Accessed: 2025-09-04.
- Spelmezan et al. (2009) Daniel Spelmezan, Mareike Jacobs, Anke Hilgers, and Jan Borchers. 2009. Tactile motion instructions for physical activities. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Boston, MA, USA) (CHI ’09). Association for Computing Machinery, New York, NY, USA, 2243–2252. doi:10.1145/1518701.1519044
- Stawarz et al. (2025) Katarzyna Stawarz, Aluna Everitt, and Judy Bowen. 2025. Exploring Opportunities for Flexible Wearables to Support Physical Training. In Proceedings of the 2025 ACM Designing Interactive Systems Conference (DIS ’25). Association for Computing Machinery, New York, NY, USA, 2156–2170. doi:10.1145/3715336.3735703
- Tajadura-Jimenez et al. (2022) Ana Tajadura-Jimenez, Judith Ley-Flores, Omar Valdiviezo, Aneesha Singh, Milagrosa Sanchez-Martin, Joaquin Diaz Duran, and Elena Márquez Segura. 2022. Exploring the Design Space for Body Transformation Wearables to Support Physical Activity through Sensitizing and Bodystorming. In Proceedings of the 8th International Conference on Movement and Computing (Chicago, IL, USA) (MOCO ’22). Association for Computing Machinery, New York, NY, USA, Article 23, 9 pages. doi:10.1145/3537972.3538001
- Takahashi et al. (2024) Akifumi Takahashi, Yudai Tanaka, Archit Tamhane, Alan Shen, Shan-Yuan Teng, and Pedro Lopes. 2024. Can a Smartwatch Move Your Fingers? Compact and Practical Electrical Muscle Stimulation in a Smartwatch. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 2, 15 pages. doi:10.1145/3654777.3676373
- Tennent et al. (2020) Paul Tennent, Joe Marshall, Vasiliki Tsaknaki, Charles Windlin, Kristina Höök, and Miquel Alfaras. 2020. Soma Design and Sensory Misalignment. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3313831.3376812
- Tessmer et al. (2022) Lavender Tessmer, Ganit Goldstein, Guillermo Herrera-Arcos, Volodymyr Korolovych, Rachel Bellisle, Cody Paige, Christopher Shallal, Atharva Sahasrabudhe, Hugh Herr, Svetlana Boriskina, Dava Newman, and Skylar Tibbits. 2022. 3D Knit Spacesuit Sleeve with Multifunctional Fibers and Tunable Compression.
- Tsai et al. (2022) Hsin-Ruey Tsai, Yu-So Liao, and Chieh Tsai. 2022. ImpactVest: Rendering Spatio-Temporal Multilevel Impact Force Feedback on Body in VR. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 356, 11 pages. doi:10.1145/3491102.3501971
- Turmo Vidal et al. (2023) Laia Turmo Vidal, Yinchu Li, Martin Stojanov, Karin B Johansson, Beatrice Tylstedt, and Lina Eklund. 2023. Towards Advancing Body Maps as Research Tool for Interaction Design. In Proceedings of the Seventeenth International Conference on Tangible, Embedded, and Embodied Interaction (Warsaw, Poland) (TEI ’23). Association for Computing Machinery, New York, NY, USA, Article 20, 14 pages. doi:10.1145/3569009.3573838
- Turmo Vidal et al. (2025) Laia Turmo Vidal, Annika Waern, Rosa Cabanas-Valdés, Lauren van Loo, Yinchu Li, and Karthik Venkataraman Meenaakshisundaram. 2025. Towards Personalized Physiotherapy through Interactive Machine Learning: A Conceptual Infrastructure Design for In-Clinic and Out-of-Clinic Support. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 670, 19 pages. doi:10.1145/3706598.3713823
- Turmo Vidal et al. (2020) Laia Turmo Vidal, Hui Zhu, and Abraham Riego-Delgado. 2020. BodyLights: Open-Ended Augmented Feedback to Support Training Towards a Correct Exercise Execution. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. doi:10.1145/3313831.3376268
- Turmo Vidal et al. (2021) Laia Turmo Vidal, Hui Zhu, Annika Waern, and Elena Márquez Segura. 2021. The Design Space of Wearables for Sports and Fitness Practices. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 267, 14 pages. doi:10.1145/3411764.3445700
- Velloso et al. (2013) Eduardo Velloso, Andreas Bulling, and Hans Gellersen. 2013. MotionMA: motion modelling and analysis by demonstration. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Paris, France) (CHI ’13). Association for Computing Machinery, New York, NY, USA, 1309–1318. doi:10.1145/2470654.2466171
- Wang and Li (2023) Jun Wang and Suyi Li. 2023. Building intelligence in the mechanical domain—Harvesting the reservoir computing power in origami to achieve information perception tasks. Advanced Intelligent Systems 5, 9 (2023), 2300086.
- Wang et al. (2016) Qi Wang, Marina Toeters, Wei Chen, Annick Timmermans, and Panos Markopoulos. 2016. Zishi: A Smart Garment for Posture Monitoring. In Proceedings of the 2016 CHI Conference Extended Abstracts on Human Factors in Computing Systems (San Jose, California, USA) (CHI EA ’16). Association for Computing Machinery, New York, NY, USA, 3792–3795. doi:10.1145/2851581.2890262
- Weng et al. (2025) Jian-Jia Weng, Calvin Ku, Jo Chien Wang, Chih-Jen Cheng, Tica Lin, Yu-An Su, Tsung-Hsun Tsai, You-Yi Lin, Lun-Wei Ku, Hung-Kuo Chu, and Min-Chun Hu. 2025. Bridging Coaching Knowledge and AI Feedback to Enhance Motor Learning in Basketball Shooting Mechanics Through a Knowledge-Based SOP Framework. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 989, 20 pages. doi:10.1145/3706598.3713324
- Wittchen et al. (2025) Dennis Wittchen, Alexander Ramian, Nihar Sabnis, Richard Böhme, Christopher Chlebowski, Georg Freitag, Bruno Fruchard, and Donald Degraen. 2025. CollabJam: Studying Collaborative Haptic Experience Design for On-Body Vibrotactile Patterns. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 1132, 20 pages. doi:10.1145/3706598.3713469
- Woźniak et al. (2020) Mikołaj P. Woźniak, Julia Dominiak, Michał Pieprzowski, Piotr Ładoński, Krzysztof Grudzień, Lars Lischke, Andrzej Romanowski, and Paweł W. Woźniak. 2020. Subtletee: Augmenting Posture Awareness for Beginner Golfers. Proc. ACM Hum.-Comput. Interact. 4, ISS, Article 204 (Nov. 2020), 24 pages. doi:10.1145/3427332
- Wu et al. (2024) Di Wu, Emily Guan, Yunjia Zhang, Hsuanju Lai, Qiuyu Lu, and Lining Yao. 2024. Waxpaper Actuator: Sequentially and Conditionally Programmable Wax Paper for Morphing Interfaces. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 865, 16 pages. doi:10.1145/3613904.3642373
- Xu et al. (2025) Ziao Xu, Chentian Zhang, Faqiang Wang, Jianyong Yu, Gang Yang, Roman A Surmenev, Zhaoling Li, and Bin Ding. 2025. Smart textiles for personalized sports and healthcare. Nano-Micro Letters 17, 1 (2025), 1–39.
- Yang et al. (2024) Yue Yang, Lei Ren, Chuang Chen, Bin Hu, Zhuoyi Zhang, Xinyan Li, Yanchen Shen, Kuangqi Zhu, Junzhe Ji, Yuyang Zhang, Yongbo Ni, Jiayi Wu, Qi Wang, Jiang Wu, Lingyun Sun, Ye Tao, and Guanyun Wang. 2024. SnapInflatables: Designing Inflatables with Snap-through Instability for Responsive Interaction. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 342, 15 pages. doi:10.1145/3613904.3642933
- Yang et al. (2023) Yue Yang, Lei Ren, Chuang Chen, Xinyue Wang, Yitao Fan, Yilin Shao, Kuangqi Zhu, Jiaji Li, Qi Wang, Lingyun Sun, Ye Tao, and Guanyun Wang. 2023. E-Orthosis: Augmenting Off-the-Shelf Orthoses with Electronics. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 625, 15 pages. doi:10.1145/3544548.3581471
- Young et al. (2019) Eric M. Young, Amirhossein H. Memar, Priyanshu Agarwal, and Nick Colonnese. 2019. Bellowband: A Pneumatic Wristband for Delivering Local Pressure and Vibration. In 2019 IEEE World Haptics Conference (WHC). 55–60. doi:10.1109/WHC.2019.8816075
- Yu et al. (2024) Xingyao Yu, Benjamin Lee, and Michael Sedlmair. 2024. Design Space of Visual Feedforward And Corrective Feedback in XR-Based Motion Guidance Systems. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 723, 15 pages. doi:10.1145/3613904.3642143
- Zhang et al. (2024) Zining Zhang, Jiasheng Li, Zeyu Yan, Jun Nishida, and Huaishu Peng. 2024. JetUnit: Rendering Diverse Force Feedback in Virtual Reality Using Water Jets. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 136, 15 pages. doi:10.1145/3654777.3676440
- Zhou et al. (2023) Qiushi Zhou, Louise Grebel, Andrew Irlitti, Julie Ann Minaai, Jorge Goncalves, and Eduardo Velloso. 2023. Here and Now: Creating Improvisational Dance Movements with a Mixed Reality Mirror. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 183, 16 pages. doi:10.1145/3544548.3580666
- Zhu et al. (2022) Junyi Zhu, Yuxuan Lei, Aashini Shah, Gila Schein, Hamid Ghaednia, Joseph Schwab, Casper Harteveld, and Stefanie Mueller. 2022. MuscleRehab: Improving Unsupervised Physical Rehabilitation by Monitoring and Visualizing Muscle Engagement. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology (Bend, OR, USA) (UIST ’22). Association for Computing Machinery, New York, NY, USA, Article 33, 14 pages. doi:10.1145/3526113.3545705