### v2

# FlexGuard: A Design Space for On-Body Feedback for Safety Scaffolding in Strength Training

Abstract.
Strength training carries inherent safety risks when exercises are performed without supervision. While haptics research has advanced, there remains a gap in how to integrate on-body feedback into intelligent wearables. Developing such a design space requires experiencing feedback in context, yet obtaining functional systems is costly. By addressing these challenges, we introduce FlexGuard, a design space for on-body feedback that scaffolds safety during strength training. The design space was derived from nine co-design workshops, where novice trainees and expert trainers DIY’d low-fidelity on-body feedback systems, tried them immediately, and surfaced needs and challenges encountered in real exercising contexts. We then evaluated the design space through speed dating, using storyboards to cover the design dimensions. We followed up with workshops to further validate selected dimensions in practice through a proof-of-concept wearable system prototype, examining how on-body feedback scaffolds safety during exercise. Our findings extend the design space for sports and fitness wearables in the context of strength training.
1. INTRODUCTION
Strength training is a widely practiced method for improving strength, mobility, and overall health. However, errors such as joint misalignment or overextension can increase the risk of injury, especially for novice practitioners (Fisher, 2025). Personal trainers help mitigate these risks by providing tailored feedback that cultivates trainees’ awareness of potentially unsafe movement and offers moment-to-moment physical support during execution. Such in-person supervision, however, is costly and depends on others’ availability, making it difficult to scale. As a result, there is a need for alternative forms of support that can scaffold trainees’ experience by helping them remain aware of safety-relevant movement while also providing embodied assistance during training. Advances in haptics now enable the emulation of physical sensations commonly used in in-person coaching, such as taps, pressure, and resistance. This creates an opportunity for intelligent wearable systems that can deliver on-body feedback to scaffold safety during strength training. What remains missing, however, is the design knowledge needed to integrate such on-body feedback into wearable systems.
Designing on-body feedback systems is challenging because users must physically experience feedback during strength training (Tennent et al., 2020; Ley-Flores et al., 2024), and obtaining functional systems is costly (Caserman et al., 2021). Investing heavily in a prototype to serve as a technology probe (Hutchinson et al., 2003) is often impractical in early-stage exploration. Moreover, any probing system must be representative of the diverse on-body feedback mechanisms that have been studied and developed. Together, these challenges hinder systematic investigation of the design space.
By addressing these challenges, we developed FlexGuard, a design space for on-body feedback to scaffold safety in strength training. We began by surveying the landscape of feedback modalities, including vibrotactile (Huang et al., 2025; Elsayed et al., 2020), pneumatic (Delazio et al., 2018; Pardomuan et al., 2024), and other approaches (Kwon et al., 2022; Lin et al., 2025b), and grouped them into two categories: extrinsic, which require active actuation (e.g., vibration or air pressure), and intrinsic, which rely on material properties (e.g., orthoses or braces) to provide passive feedback once configured. We then prepared technology probes representing these categories, along with supporting props, and conducted nine co-design workshops in our institution’s gym. Each session paired a novice trainee with an expert trainer, who collaboratively DIY’d personalized low-fidelity on-body feedback systems for five representative exercises and tested them immediately in practice. Through thematic analysis, we distilled the needs and challenges that informed the derivation of FlexGuard.
FlexGuard is organized around two complementary dimensions: Sensing & Triggering and Feedback Delivery. Sensing & Triggering specifies which indicators should be monitored (e.g., joint angle, movement dynamics, motion path, muscle activation, or inter-limb coordination), how triggering policies are defined (e.g., expert-based or personalized), and where monitoring should occur (e.g., target, stabilizing, moving, or compensation regions). Feedback Delivery describes how feedback is provided once triggered, including the body zones where different feedback types are applied, the intensity of feedback, its directional strategies, and its timing across movement phases. Together, these dimensions define how on-body feedback can be timed, targeted, and configured to scaffold safety during exercise.
We ran three speed dating workshops with seven trainees with limited strength training experience. Participants first performed exercises with the provided equipment in our lab to ground their reflections and then discussed a series of storyboard scenarios that varied Sensing & Triggering and Feedback Delivery, spanning the dimensions of FlexGuard. The results indicated that participants found the speculative scenarios relevant and useful. These findings provided conceptual validation of FlexGuard and allowed us to refine the design space based on evidence of user alignment with its dimensions. n addition, we identified underlying needs across scenarios, including expectations around feedback semantics, timing and progressivity, and adaptability across users, exercises, and contexts.
We further validated selected dimensions of FlexGuard in practice by developing a proof-of-concept (PoC) wearable system. Motivated by the speed dating findings, we narrowed the scope of the PoC to focus on intrinsic feedback and the underlying user needs identified earlier. As wearable systems move toward personalized intelligent support, intrinsic feedback becomes particularly relevant due to its potential for continuous and immediate embodied assistance. However, prior work provides limited design knowledge on how intrinsic mechanisms should be integrated with real movement contexts. To address this gap, we implemented a strap-based intrinsic-feedback suit tested on two exercises that together span two primary degrees of freedom of the shoulder. A camera-based system detected deviations from target posture and signaled strap adjustments, which were manually enacted in this PoC to simulate an automated system. This setup allowed participants to experience phase-specific variations in feedback direction and intensity, providing embodied validation of the design space and confirming key assumptions from the speed dating study in practice.
Our contributions are as follows:
We introduce FlexGuard, a design space for on-body feedback that scaffolds safety in strength training, grounded in nine co-design workshops with trainers and trainees.
We articulate dimensions that characterize how systems can sense and deliver feedback to scaffold trainees’ awareness of and support for safety-relevant movement during exercise.
We present two complementary evaluations: a speed dating workshop and a PoC wearable system, together demonstrating how FlexGuard can guide the design of adaptable on-body feedback systems.
2. RELATED WORK
Our research builds upon prior work in four key areas: Safety and Risk in Strength Training, Sensing and Feedback Systems for Physical Training, Advances in Materials and Wearable Technologies, Intelligent Systems in Movement and Skill Learning.
2.1. Safety and Risk in Strength Training
Strength training carries inherent safety considerations, particularly when exercises involve high external loads that stress the musculoskeletal system. Body regions that both bear the load and move dynamically, such as the lower back and shoulders, are especially prone to overload or misalignment-related issues (Hedrick and Wada, 2008; Barnes et al., 2021). While such issues are common, they are often avoidable. Research indicates that a large portion of lower back strain incidents are linked to improper lifting form, excessive loads, and overtraining, which all are factors that can be mitigated through proper technique and training practices (Alqarni, 2019). Faulty alignment and compensatory movement patterns introduce abnormal stress across joints and soft tissues, contributing not only to unsafe strain but also to long-term issues. This risk is especially pronounced in novice trainees, whose motor control systems may not yet be developed enough to coordinate complex compound movements effectively (Aslam et al., 2025). Thus, consistent attention to motor learning is fundamental for safe and effective strength training (Latella et al., 2019).
This is why working with a personal trainer is widely recommended, particularly for beginners. Professional supervision ensures that trainees are taught correct movement techniques, are properly spotted to maximize repetitions safely, and are guided through appropriate loading progressions. When properly coached, strength training is a safe activity relative to other sports (Fisher et al., 2023). In contrast, unsupervised training is associated with higher risks due to inconsistent technique and overly aggressive load selection (Faigenbaum and Myer, 2010; Gabbett, 2020). While technologies exist to support self-coaching (Snouzy, 2025), human coaches provide multifaceted guidance that extends beyond form correction and motor habit development to include enforcing safety practices such as spotting, load monitoring, and early detection of execution breakdowns (Fisher et al., 2023).
In addition to coaching, trainees frequently use supportive wearable gear to enhance safe training practices. Equipment such as weightlifting belts, wrist wraps, knee sleeves, and elbow sleeves is commonly used in gyms to help maintain alignment and distribute stress during heavy lifts. Weightlifting belts, for instance, are used to increase intra-abdominal pressure and reduce spinal loading during squats and deadlifts (Herbaut and Tuloup, 2025). Knee sleeves improve proprioception during deep knee flexion movements (Machek et al., 2021). Users can also define and produce custom wearables for their specific needs (Xu et al., 2025). For example, tailoring their own knee patch or squat suit to support heavy lifting (Latella et al., 2024).
FlexGuard builds on this trajectory by extending safety considerations beyond passive support and human supervision. Leveraging haptics for on-body sensation, FlexGuard advances toward adaptive systems that actively sense and respond to cues and deviations during exercise, providing intuitive feedback to support safety scaffolding during training.
2.2. Sensing and Feedback Systems for Physical Training
Designing wearables requires careful consideration of how the body perceives and responds to interaction, emphasizing not just function but felt experience (Tennent et al., 2020; Windlin et al., 2022); thus, there are two key components: sensing and feedback methods on the body. This aligns with the principles of soma-based design theory, which foregrounds the subjective, lived body as a central resource for interaction design (Höök et al., 2017). Ley-Flores et al. used a ‘somatic dress-up’ approach to explore how negative body perceptions can hinder engagement in physical activity (Ley-Flores et al., 2024). Sensing technologies have enabled researchers and designers to monitor a wide range of body states during physical training. Researchers have explored the use of sensors to monitor muscle activity (Zhu et al., 2022; Esposito et al., 2018; Rudolph et al., 2022). Systems like RecoFit (Morris et al., 2014) and LiftSmart (Kowsar et al., 2019) use wearable sensors to monitor and assess users’ exercise performance. ProxiFit (Kim et al., 2023) senses exercise repetitions by detecting magnetic variations during movement. Other systems rely on pressure or tactile sensors to provide posture-related sensing, such as GymSoles, which uses insole-based pressure readings to visualize a user’s center of pressure during lifts (Elvitigala et al., 2019). FormFit (Rao et al., 2024) employs computer vision for automatic pose correction, while Zishi (Wang et al., 2016) combines smart textiles and motion sensing to monitor posture during rehabilitation. More integrated systems like E-Orthosis (Yang et al., 2023) embed sensing directly into orthotic devices using textile electrodes, enabling real-time feedback and data logging in rehabilitation and sports contexts. Researchers have also studied textile-based knee strain sensors for real-time monitoring during workouts (Aliyana et al., 2024), as well as developable shoulder sensors for precise movement tracking (Evke et al., 2019). Commercial platforms such as Catapult (Catapult, 2025) offer elite-level sports analytics, although they often remain inaccessible or inflexible for everyday users.
A parallel stream of work focuses on feedback modalities for physical training. Researchers have explored output devices such as smart rings (Wang et al., 2025), along with visual augmentation via projection and AR/VR (Yu et al., 2024). Examples include systems like BodyLights (Turmo Vidal et al., 2020), augmented instructional videos (Semeraro and Turmo Vidal, 2022), real-time overlays (Keelawat and Suzuki, 2024; Ma et al., 2024), and spatial visualizations in FlowAR (Jo et al., 2023a), all of which guide users through contextual cues during training. Some systems, like Weight-Mate (Paay et al., 2020), combine visual and auditory feedback to help users maintain alignment. Haptic feedback systems have also gained attention, including vibration-based alerts for physical activities (Spelmezan et al., 2009; Kosmalla et al., 2016), implemented in TrainerTap (Jo et al., 2023b), VibraForge (Huang et al., 2025) and Subtletee (Woźniak et al., 2020). Such systems may be particularly beneficial for individuals with visual impairments when practicing yoga (Islam et al., 2022). Han et al. (Han et al., 2023) explore 3D-printed tactor geometries for generating diverse haptic sensations. EMS-based stimulation for movement correction is another promising direction. For instance, FootStriker helps improve running foot form to prevent knee-related injuries (Hassan et al., 2017). Additionally, the Teslasuit (Caserman et al., 2021) exemplifies a production-grade EMS and TENS platform capable of training full-body motion. Wearable robotics such as HapticSnakes, which use waist-worn robotic arms to deliver multi-haptic feedback (including taps and brushing) to various locations across the upper and lower body (Al-Sada et al., 2020). Other systems explore novel modalities, including skin-stretching for proprioceptive illusions (Lin et al., 2024b), pneumatic water-jet feedback (Zhang et al., 2024), localized pressure for knee joint stabilization (Park et al., 2020), and torque-based force guidance for arm motion (Kim and Asbeck, 2022). While significant work has focused on the upper body, some systems involve skin-stretching on the legs to provide proprioceptive feedback for the lower body, demonstrating the applicability of on-body cues to other regions (Wang et al., 2020).
Researchers have studied broader challenges in the context of sports technologies (Turmo Vidal et al., 2021; Elvitigala et al., 2024), highlighting the need for systems that can adapt to individual physical contexts and training goals. Personalization is especially emphasized in recent work on sports wearables (Stawarz et al., 2025), while others call for more interpretability and objective relevance in how feedback is delivered (Tajadura-Jimenez et al., 2022).
These works show the breadth of sensing and feedback modalities available, but most rely on predefined mappings and offer limited user configurability. Even state-of-the-art products like the Teslasuit (Caserman et al., 2021) require heavy reliance on expert setup. FlexGuard addresses this gap by enabling end users to configure, time, and adjust feedback delivery interactively, while maintaining safety by anchoring its logic in expert knowledge.
2.3. Advances in Materials and Wearable Technologies
Recent developments in soft robotics, programmable materials, and digital fabrication have expanded the design space for wearable systems. Pneumatic actuation systems like PneumAct (Günther et al., 2019), VabricBeads (Pardomuan et al., 2024), Force Jacket (Delazio et al., 2018), and Frozen Suit (Al Maimani and Roudaut, 2017) use air or liquid pressure to generate movement, resistance, or stiffness changes in the body. milliMorph (Lu et al., 2019) pushes this further by introducing millimeter-scale fluidic chambers for fine-grained, fluid-driven shape change. PneuMa (Saini et al., 2024) introduces pneumatic bodily extensions to support daily movement. Actuation mechanisms inspired by biological forms, such as artificial muscles (Li et al., 2017), offer compact and powerful actuation possibilities. Researchers have also investigated collaborative approaches for designing wearables with on-body feedback such as vibrotactile patterns (Wittchen et al., 2025; Sung et al., 2024; Zhou et al., 2025). Furthermore, to enhance further capability, tools like PneuMod enable computational design and fabrication of pneumatic actuators with programmable material properties, allowing for fine-tuned force and stiffness responses (Zhang and Sra, 2021).
Programmable materials with bistable or multistable properties allow for richer forms of mechanical feedback and passive guidance (Alu et al., 2025; Lee et al., 2026). Works such as Kirigami Haptic Swatches (Chang et al., 2020), FlexTure (Johnson et al., 2023), Waxpaper Actuator (Wu et al., 2024), and ConeAct (Lin et al., 2024a) explore structural snapping transitions, geometric locking, and reactive haptics. Snap-through mechanisms, like those in SnapInflatables (Yang et al., 2024), enable rapid shape changes with localized feedback. Springlets (Hamdan et al., 2019) introduces shape-memory alloy springs that simulate on-skin sensations such as touching, pinching, pulling, etc.
Wearable fabric-based sensing and actuation is also advancing rapidly (Li and Luo, 2024). SeamFit introduces garments capable of logging exercise movements (Yu et al., 2025). Luo et al. demonstrate textile-based control for skill transfer tasks like piano playing (Luo et al., 2024), while ReKnit-Care (Hu et al., 2025) combines knitted sensors with EMS to provide fine-grained hand feedback for rehabilitation. Lin et al. (Lin et al., 2025a) introduce passive microstructures to switch between states while worn to change perceived physical environment. Tessmer et al. designed a spacesuit with tunable compression (Tessmer et al., 2022). FlexGuard draws from this material ecosystem, integrating advances in actuation and textiles to explore how adaptive, on-body feedback can support safety scaffolding and personalization during strength training.
2.4. Intelligent Systems in Movement and Skill Learning
Interactive Machine Learning (IML) refers to systems that allow end-users to guide, correct, or shape the learning process through iterative interaction. Unlike traditional supervised ML, IML invites user-in-the-loop contributions during model refinement (Amershi et al., 2014). Hartmann et al. (Hartmann et al., 2007) emphasize the value of interactivity in enabling more intuitive and user-aligned learning. These systems combine sensor data, user feedback, and pattern recognition to empower non-expert users to build and refine interactive logic.
In movement learning and demonstration-based systems, researchers have applied IML techniques to encode and reuse human motion, which often involves tacit knowledge that is difficult to articulate explicitly (Gillies, 2019). MotionMA (Velloso et al., 2013) enables users to demonstrate movement patterns that the system can model and compare in real time. Mimic (Mahadevan et al., 2022) applies similar principles in teleoperation, letting users record and re-trigger physical behaviors. Other approaches like danceON (Payne et al., 2021) and Rapido (Leiva et al., 2021) focus on embodied programming, through specification or demonstration, for creative and spatial applications. Vidal et al. propose a framework and infrastructure for applying IML in physiotherapy to support both in-clinic and out-of-clinic patients (Turmo Vidal et al., 2025). Haptic Scores (Escudero, 2024) attempts to map detected movements to embodied haptics for dance with EMS feedback. In sport-specific contexts, Weng et al. decompose movement into feedback-friendly units to integrate AI-driven textual coaching for basketball shooting (Weng et al., 2025). However, such textual feedback approaches (Mahadevan et al., 2024; Pu et al., 2025) may be less effective in physical tasks like strength training. FlexGuard builds on these directions by embedding interactive intelligence directly into the wearable interface, enabling trainees to shape feedback for safer and more personalized strength training.
3. LANDSCAPE OF FEEDBACK MODALITIES
Wearable systems have explored a wide spectrum of on-body feedback modalities to support physical interaction and learning. Our review of over 50 works from wearable computing, ubiquitous computing, and haptics research, and relevant examples from gym training practices and soft robotics surfaces several recurring strategies: vibration, electrical stimulation, pneumatic actuation, thermal cues, mechanical-based feedback, and structural supports. Table 1 summarizes representative systems and approaches across these categories.
| Modality | Examples | Description |
| Vibration | TrainerTap (Jo et al., 2023b), VibraForge (Huang et al., 2025), HapticHead (Kaul and Rohs, 2017), Zishi (Wang et al., 2016), HapticPilot (Sung et al., 2024) | Pulses or patterns on the skin. |
| Electrical (EMS / TENS) | Teslasuit (Caserman et al., 2021), FootStriker (Hassan et al., 2017), Smartwatch EMS (Takahashi et al., 2024), ReKnit-Care (Hu et al., 2025), Haptic Scores (Escudero, 2024) | Electrical stimulation of muscles or nerves. |
| Pneumatic | VabricBeads (Pardomuan et al., 2024), PneumAct (Günther et al., 2019), Force Jacket (Delazio et al., 2018), FrozenSuit (Al Maimani and Roudaut, 2017), aSpire (Choi et al., 2022), JetUnit (Zhang et al., 2024), OmniFiber (Kilic Afsar et al., 2021) | Air or fluid pressure for force or resistance. |
| Thermal | ThermoVR (Peiris et al., 2017), Augmented Breathing (Brooks et al., 2024) | Heating or cooling sensations. |
| Mechanical | Springlets (Hamdan et al., 2019), ArmDeformation (Lin et al., 2024b), ImpactVest (Tsai et al., 2022) | Mechanically generated forces or deformations. |
| Structural Supports | Bistable Orthoses (Lin et al., 2025b), Stiffening Garment (Kwon et al., 2022), SkinMorph (Kao et al., 2018), ThermoFit (Wang et al., 2023), gym gear (e.g., knee sleeves (Machek et al., 2021), squat suit (Latella et al., 2024), lifting belts (Herbaut and Tuloup, 2025)) | Passive supports for posture or training. |
Across these categories, we distinguish between extrinsic mechanisms (externally actuated by components outside the wearable material, such as vibration, electrical, pneumatic, thermal, or material-driven actuation that deliver forces or sensations to the body) and intrinsic mechanisms (intrinsically produced by the wearable’s material itself; passive supports such as braces, sleeves, or garments that guide or resist movement through direct material contact). Since strength training demands both corrective cues and physical support, our subsequent sections of this paper focus on these two modalities through co-design workshops exploring their affordances in training contexts.
4. ON-BODY FEEDBACK FOR STRENGTH TRAINING
To understand the needs and challenges of on-body feedback systems for strength training, we conducted a co-design workshop study with novice trainees, supported and informed by expert trainers. To achieve that goal, we have to prepare materials for the study including technology probe and worksheets, that will cover comprehensive space of strength training.
We focused on a set of upper-body strength exercises where errors could compromise safety, to scope the study. Although not full-body, the selected exercises encompass a broad range of movement demands in strength training, offering insights that can largely generalize to other body regions and exercises. The representative set included five exercises: dumbbell front hold, dumbbell incline fly, bench press, lateral raise, and Arnold press. Each exercise highlighted a distinct movement error risk, such as posture holding, range of motion (ROM), inter-limb balance, tempo, or complex transitions. Further details about safety-relevant factors are shown in Figure 10.
4.1. Technology Probe Design for DIY Low-Fi Prototyping
To enable rapid DIY prototyping and immediate exploration of feedback sensations, we developed a technology probe (Hutchinson et al., 2003; Ley-Flores et al., 2024) designed for quick reconfiguration. The base system included adjustable shoulder braces and elbow sleeves, equipped with customizable feedback components such as inflatable air wedges on the shoulders, and rigid and stretchy straps on the elbows. The probe supported two distinct modes of on-body feedback:
4.1.1. Extrinsic Feedback
Externally actuated feedback outside of the wearable’s material structure, simulated through the modulation of air wedges.
4.1.2. Intrinsic Feedback
Intrinsically generated feedback arising from the wearable’s own materials, achieved through tension in rigid or elastic straps that respond directly to the user’s movement.
We selected these modalities for the study based on their practicality for rapid prototyping in a gym environment (e.g., without electronic wiring) and their coverage of diverse experiences. The air wedges, serving as extrinsic feedback, allowed participants to explore a range of sensations on demand by varying pumping patterns and intensity. In contrast, intrinsic feedback depended almost entirely on material properties, so we provided both rigid and elastic options to offer participants a broader range of experiences.
Velcro interfaces allowed participants to quickly attach, reposition, or remove components as needed. To further support collaborative prototyping, we provided additional materials such as duct tape, resistance bands, and a mini-mannequin to facilitate communication between trainers and trainees. This setup enabled hands-on experimentation: participants configured feedback on their own bodies, performed representative movements, and reflected on the sensations in real time (Figure 2).
4.2. Worksheet Design
We prepared a worksheet for each exercise. Each worksheet included body maps (Turmo Vidal et al., 2023) with figures of a person performing the exercise at representative moments, such as the starting position, mid-movement, and the maximum extension point. For the dumbbell front hold, an additional figure depicted an isometric hold at 15 seconds. These figures served as reference frames for participants to annotate, with the option of requesting extra sheets to sketch alternative designs or capture other parts of the movement cycle. As body maps are adaptable to diverse research contexts (Turmo Vidal et al., 2023), in our study they functioned both as a tool to guide the session and as an open-ended space for dialogue between trainer and trainee. They were designed for participants to sketch low-fidelity prototype designs, align on preferred configurations, and highlight specific body regions requiring attention. An example worksheet can be seen in Figure 3.
4.3. Study Protocol
We conducted nine co-design sessions, each involving a matched trainer–trainee pair. Table 2 summarizes participant demographics, training experience, and notable athletic backgrounds.
| Session | Trainer (Age, Gender) | Strength Training Experience | People Coached | Other Sports & Fitness Background | Trainee (Age, Gender) |
| S1 | E1 (48, M) | 2 yrs | 5+ | – | N1 (23, M) |
| S2 | E2 (26, M) | 1.5 yrs | 5–10 | – | N2 (26, M) |
| S3 | E3 (25, F) | 1 yr | 2 | – | N3 (33, F) |
| S4 | E4 (28, M) | 1 yr | 2 | – | N4 (24, M) |
| S5 | E5 (24, F) | 8 yrs | 5+ | • Licensed Zumba & Aerobics instructor • Experience teaching classes with handheld weights | N5 (26, F) |
| S6 | E6 (23, F) | 9 yrs | 5+ | • Division I track/cross-country athlete • Provided instruction on form and spotting to peers and juniors | N6 (23, F) |
| S7 | E7 (24, F) | 5-7 yrs | 20+ | • National-level rowing athlete • Instructor of rowing camps | N7 (33, F) |
| S8 | E8 (22, M) | 8 yrs | 100+ | • Fitness instructor (3 years) • Track/cross-country athlete: PAC-12 Honor Roll, National U-20 champion | N8 (25, M) |
| S9 | E9 (25, M) | 8 yrs | 30+ | • Gym trainer (5 months) • Trained peers in calisthenics club | N9 (31, M) |
All novice trainees (N1–N9) had less than three months of consistent strength training experience, while expert trainers (E1–E9) were recruited based on having at least one year of consistent training experience and prior coaching experience to ensure that trainees received appropriate guidance during the study.
Participants were instructed to wear shirts with sleeves, as the technology probe was designed to be worn on the shoulders and arms. To ensure hygiene and minimize direct skin contact, we provided disposable elbow sleeves for participants to wear under the probe components.
Sessions took place in a private studio room at our institution’s gym and lasted approximately 90 minutes. To ensure participant comfort, especially during on-body probe configuration, each pair was matched by gender. All participants provided informed consent for video and audio recording. When participants used public gym equipment (e.g., for the bench press), continuous audio recording was maintained, and short video clips were recorded when appropriate, focusing only on consenting individuals. Throughout the session, participants were encouraged to think aloud and take breaks as needed. Each participant received $37.50 in compensation ($25 per hour) for their time and effort.
Each session followed the structure below:
4.3.1. Initial Interview
Each session began with two brief, separate audio-recorded interviews, one with the trainee and one with the trainer. Participants discussed their individual training backgrounds, any prior injuries, experiences, or safety concerns related to strength training, and their experience coaching or being coached. These interviews provided context for interpreting their design decisions during the rest of the session.
4.3.2. Worksheet Introduction
We introduced the study using worksheets to guide participants through the activities. We explained the project and highlighted three design goals to optimize:
DG1: Promote safety or prevent injuries by providing feedback.
DG2: Deliver on-body sensations.
DG3: Personalize feedback to the individual needs of each trainee.
4.3.3. Probe Demonstration
Next, the research team introduced the technology probe. We first demonstrated the default configuration of the probe (Figure 2A). Participants were then encouraged to experiment with customizing the setup to their own preferences, such as attaching straps in alternative positions (e.g., waist-to-shoulder). Both trainer and trainee wore the probe and explored three core feedback configurations:
Inflatable air wedges (extrinsic, active pressure feedback)
Rigid straps (intrinsic, non-elastic passive resistance)
Stretchy straps (intrinsic, elastic passive resistance)
This phase allowed participants to experience the tactile qualities of each configuration and develop an on-body sense of how feedback might support training. Since our focus was on personalizing feedback for trainees (DG3), trainers were not required to wear the probe beyond this stage.
4.3.4. Typical Training without Feedback
To establish a baseline and observe natural trainer–trainee dynamics, the trainee removed or loosened the probe, and the trainer guided them through a typical coaching session using minimal weight. Heavier weights were available to explore more challenging scenarios, but their use was optional, supervised by the trainer, and contingent on the trainee’s comfort and consent. Participants were encouraged to take breaks or drink water during the study. This phase helped surface real-world challenges, and implicit feedback strategies.
4.3.5. Co-Design Activity and Reflection
After the initial training, each trainer–trainee pair engaged in a collaborative co-design activity by sketching their ideas on the worksheet. They then configured the probe to reflect these design decisions, after which the trainee re-performed the exercise while wearing the configured probe. Participants reflected on how the feedback felt, discussed potential improvements, and recorded notes on the worksheet. If time permitted, they iterated on the design to better align with their intentions. The overall DIY process is demonstrated in Figure 4
4.4. Data Analysis
We conducted a thematic analysis using reflexive approaches that combined both deductive and inductive coding (Braun and Clarke, 2019) of video and audio recordings as well as annotated worksheets. The first author developed initial codes informed by theoretical frameworks from prior review articles on sports wearable technologies (e.g., (Turmo Vidal et al., 2021; Seçkin et al., 2023)), focusing on themes such as inputs, outputs, and their interplay. Examples of early codes included “joint deviation cue” (input), “vibrotactile” (output), and “synchronous execution” (interplay), etc. After collaboratively reviewing two sessions with the fourth and seventh authors, including videos (Knoblauch et al., 2014), transcripts, and worksheets, the first author generated preliminary groupings, which were then shared with the second, third, fourth, and seventh authors to refine the emerging themes. These results were iteratively discussed and revised during team meetings with the broader author group and were also compared against more recent design space publications (e.g., (Stawarz et al., 2025)). The remaining recordings were reviewed by the fourth and seventh authors under the supervision of the first author. The resulting design space was extensively discussed and refined by the first, fourth, and seventh authors, with additional input from the last author as well as the rest of the team during meetings.
During this process, coding disagreements occasionally surfaced and were resolved through discussion. For example, when identifying the dimensions related to motion variation, one author initially preferred coding “velocity” as its own metric based on observations from lateral raise deviations, while others argued for a broader category such as “locomotion,” which could generalize to full-body exercises. Through discussion, we recognized that several correction cues extended beyond pure velocity. We therefore agreed on the more inclusive term ”movement dynamics”, which better captured the relevant changes in movement and was more appropriate for the context of this study. Another example emerged when we initially coded several dimensions using labels such as “limiting ROM extremas,” “warning,” “restricting motion path,” and “reminder.” During a team meeting, a team member noted that these labels conflated the mechanics of feedback with the user’s intended outcome. To make the dimensions agnostic to user intent, we refined them into the more neutral categories of ”opposing”, ”guiding”, ”attentional”, and ”compressive”.
5. FLEXGUARD
Drawing on insights from the co-design workshops, we introduce FlexGuard, a design space that models how on-body feedback can be sensed, triggered, and delivered to support safety scaffolding during strength training.
We structure this design space around five exercise phases: Initiation, Execution, Hold, Return, and Completion. The Execution and Return phases may be either concentric or eccentric, depending on the exercise. Dividing exercises into phases enables us to map specific design choices to context-dependent feedback opportunities.
FlexGuard comprises two overarching categories: Sensing & Triggering and Feedback Delivery. Throughout the paper, we use blue styling to denote dimensions of Sensing & Triggering, and purple styling to denote dimensions of Feedback Delivery.
5.1. Sensing & Triggering
Figure 5 summarizes the design space for Sensing & Triggering, spanning Trigger Metric, Trigger Policy, and Trigger Location. These dimensions define when and where the system should monitor to actuate feedback during strength training, as well as when to release it. Sensing must capture not only cues that warrant intervention but also signals indicating that feedback can be withdrawn once the user has corrected their form.
5.1.1. Trigger Metric
Participants identified several physical characteristics that could serve as reliable triggers for feedback. Across multiple sessions, joint angle was the most consistently mentioned marker (n=8), particularly when arms drifted out of alignment. As E3 noted, “the only thing that has to be fixed is just the angle of motion.” Trainers also cautioned against end-range errors—“People often arch their back and then tend to raise their arms. I think that might be helpful but make sure that you don’t go above range.” (S7). Others emphasized movement dynamics (e.g., acceleration or tempo) as indicators of excessive momentum rather than controlled lifting; during a lateral raise, E9 explained to N9, “Sometimes, I’m going too heavy and using too much momentum. For the last few reps, you can often try to get that extra set.” In S3, participants highlighted inter-limb coordination (e.g., symmetric pacing/height in bench press) as a cue to synchronize both sides, noted with ”It’s a problem if the right hand is further than the left hand”. In S8, feedback was tied to muscle activation, with the trainer remarking, ”…almost feel like a little bit of pressure in the spot that you’re trying to target… so you know that you’re not using the wrong muscle”. Three sessions (S2, S7, S9) discussed motion path, noting that a shaky bar or wobbling dumbbell trajectory signals fatigue and breakdown of posture. For instance, participant noted ”For this motion… If you do more up, it will hurt your joint.”
5.1.2. Trigger Policy
Participants proposed rules for deciding when feedback should be actuated and when it should be withdrawn. Two broad categories emerged. expert-based policies draw on general training knowledge, such as recommended ROM, standard joint angles, and common coaching cues. Across 6 sessions, trainers tend to transfer their expertise verbally and give interventions when trainees drifted out of widely accepted guidance. For example, most trainers emphasized enforcing limits when elbows begin to lock during bench press, ”You want to fully extend but you don’t want to lock your elbows”, or when it deviates from the shoulder-level path, ”Sometimes people will lift them like this… we don’t want that.” In contrast, across 4 sessions, participants also highlight the need for personalized policies that adapts the thresholds to individual trainee, accounting for factors such as fatigue, muscle engagement, or body type. For instance, feedback could be tuned to a user’s tolerance for fatigue or to whether the intended muscle is activating effectively, both of which can vary significantly across body types, as emphasized by E5 regarding N5’s arms length, ”Depending on how long this area is, this bubble needs to be further up as opposed to on my shoulder blade.”. While all policies, including expert-based, can ultimately be personalized, we make this distinction to separate those with only slight variations across most body types from those that are highly subjective.
5.1.3. Trigger Location
Rather than naming specific muscles, we describe locations in terms of their functional roles to allow for greater generalizability. Four categories emerged with a mean of 7 occurences by participants on each, corresponding to Figure 5. The target region was a common focus, such as in the incline dumbbell fly: “If you started overcompensating and using your wrong muscle, it would like identify that muscle and you’d know you don’t want [that]” (S6). Trainers also warned about overuse of compensation regions, where non-target areas begin to take over the movement, as noted by E6 ”If you feel an ache in another area, then you know that you’re not doing it correctly”. The importance of stabilizing regions, particularly the upper back and core, was emphasized across sessions—“If it has a baseline measurement for not engaging your core and then engaging your core, you’ll able to find your center.” (S5). Finally, participants pointed to moving regions (e.g., forearms and hands during an incline fly, or the wrists in a bench press) where feedback can guide trajectory directly. Although the muscle groups at these locations are not necessarily being worked, they act as reference points for maintaining correct movement.
5.2. Feedback Delivery
Feedback Delivery (Figure 8) captures how corrective cues are conveyed to the user once a triggering condition is met as defined in Sensing & Triggering. We define this not only as the onset of feedback, such as vibration, resistance, or pressure, but also its offset. Participants emphasized that the withdrawal of feedback was equally meaningful, since the absence of cues could serve as confirmation that posture had been corrected or movement had returned to the desired path. In this way, Feedback Delivery encompasses both the activation and the release of feedback, as well as its modulation in intensity and timing.
5.2.1. Feedback Type
We distinguish between extrinsic feedback, delivered through active actuation, and intrinsic feedback, which emerges from material properties such as straps or elastic bands.
extrinsicfeedback can be tuned to produce different actuation patterns. Participants described preferences for continuous cues that guide movement throughout a repetition—“Imagine something helping draw a line for your hands to follow” (N2) which is a preference raised in 3 out of 9 sessions—as well as discrete cues that act as one-time warnings when posture falters—“You shouldn’t nag them, just nudge once if they’re falling” (E1), emerged in 2 sessions (S1, S3).
intrinsicfeedback depends on how materials constrain or guide motion. We distinguish between rigid and elastic responses. rigid feedback provides firm constraints, such as straps that enforce strict ROM limits: “I could feel when I pulled wrong because the strap caught me” (S5). In contrast, elastic feedback allows more flexibility, offering resistance that increases with deviation and guiding users back on track without completely halting movement. As E7 explained, “So a strap that is going to help them maintain the stability. When I’m pressing it, I don’t want to go up and down… if the strap helps us to do that, then it’s pretty good.” Another participant added ”Elastic is more flexible…it wouldn’t stop me from doing the wrong movement. It would just notify me and I could readjust to my own.” Comments favoring rigid constraints appeared at 5 sessions, while preferences for elastic guidance were raised in 3 sessions.
5.2.2. Feedback Intensity
This dimension reflects how forceful the feedback feels, ranging from low to high. low intensity was favored for subtle prompts or early-stage correction (mentioned in 3 sessions). As E3 explains to N3, “When someone exercises, they will not have concentration. When it comes to shaking (haptic feedback), it’s not practical.” In contrast, high intensity was valued for safety-critical moments, such as E6’s proposal for strong restricting feedback in the return phase of a lateral raise: “If you started overcompensating and using wrong muscles, it would like identify that muscle and shift your arm back to make sure you’re not using the wrong muscle. An ache in another area would let you know that you’re not doing it correctly.” Participants also emphasized progressive intensity from at least 3 sessions (feedback that begins subtly and escalates if the error persists or worsens) as in S6 where participant highlighted ”Like slowly over time you can do less and less with the pressure.”. This approach mirrors how a trainer might tolerate small deviations but intervene more forcefully over time as errors become more critical.
5.2.3. Feedback Latency
This dimension distinguishes between instant and delayed feedback, reflecting how quickly the system should respond once a deviation is detected. Immediacy is critical in situations involving high safety risk, where feedback must intervene before damage occurs, ”If the strap helps make sure I don’t go too much in front or too much back… that prevents shoulder injury.” Four sessions emphasized the need for immediate feedback especially for shoulder and elbow. At the same time, participants recognized the value of delayed feedback that tolerates minor deviations and only escalates if errors persist. In S8, trainer states the need for guidance based on tempo not not urgent correction, ”Maybe like click click click… you’d know you’re in the correct position… help with timing and slowing movement down.” Such delayed feedback was seen as less disruptive and more supportive, as it felt less intimidating to users. Several also expressed preferences for progressive onset, where feedback begins subtly and intensifies if the deviation continues. A more tolerable or delayed feedback were discussed in three sessions. Taken together, these insights suggest that Trigger Policies must balance the safety benefits of instant correction with the patience of delayed feedback that reduces cognitive overload.
5.2.4. Feedback Location
This dimension describes where on the body feedback is delivered. From the co-design workshops, we derived body zone maps that highlight regions associated with feedback (Figure 9). extrinsic feedback can be applied directly to the same muscle group identified by Sensing & Triggering, or redirected to a different site to signal correction. In S5, participant remarks ”Bench press on wrist is the only one I think uses resistance.” A participant also annotated ”A strap in the elbow bend would help” from the co-design workshop in the dumbbell incline fly worksheet as shown in Figure 3. In contrast, intrinsic feedback is delivered at the site of motion itself, where resistance provides direct guidance. For example, constraining elbow flare or stabilizing the shoulders. Several participants highlighted how resistance at the limb could guide technique, such as S4’s observation that “If you tighten it from this point, there’s a chance that I won’t be able to drop my arm down,” underscoring how intrinsic feedback constrains movement at the source.
5.2.5. Feedback Direction
This dimension describes how corrective forces or cues are oriented relative to the body. We identified four strategies. opposing feedback was discussed in 6 of 9 sessions, to resist unwanted momentum, such as adding friction in a lateral raise to counter swinging as N3 marks ”Right now I feel the restriction in the different direction. When I try to lean my arm toward my body, I feel the restriction this way.”. 4 sessions mentioned guiding feedback nudges the user toward the correct path, for example assisting with rep completion or steering the arms back into the proper motion plane, as N6 described when correcting form during a dumbbell incline fly, ”you know, to kind of like shift your arm back to make sure you’re not using the wrong muscle.”. attentional feedback provides localized cues without displacement, such as vibration in a dumbbell front hold to warn when the arms drop below a safe angle as suggested by N1. Finally, compressive feedback applies inward pressure, for instance stabilizing the core to maintain posture, which participants in three sessions (e.g., E3, E4, E5) emphasized with their trainees, “always engage the core.”
5.3. Connecting Sensing & Feedback
The two dimensions of Sensing & Triggering and Feedback Delivery do not operate independently. Rather, they form a closed loop in which detected cues directly shape how feedback is delivered. Figure 10 illustrates this integration across five representative exercises surfaced during co-design workshops, with example approaches for addressing issues at specific Trigger Locations to scaffold safety during movement. Each row depicts an exercise across its phases of motion, highlighting the Trigger Metrics and Trigger Policies that determine when feedback is activated, alongside the corresponding feedback strategies.
This mapping underscores how expert-based policies can guide the monitoring of metrics (e.g., joint angle, motion path) at locations where safety risk is elevated. It also extends to personalized policies, such as monitoring muscle activity, which can vary across body types and help detect cases where the intended muscle is not engaged. Once a cue is flagged, Feedback Delivery defines the type and form of response. For example, in a lateral raise, exceeding a speed threshold of movement dynamics triggered intrinsic resistance to slow momentum, which could escalate with progressive intensity, from low to high intensity, if the deviation worsens.
By situating these elements side by side, the integrated design space highlights how different policies and feedback strategies can be tailored to the unique risks of each exercise. This connection makes explicit the logic of the system: cues sensed through body state and motion lead to feedback that constrains, guides, or withdraws in order to promote safe and effective strength training.
6. SPEED DATING
To evaluate the FlexGuard design space before investing in implementation, we conducted a speed dating workshop. Following prior work using this methodology (Davidoff et al., 2007; Zhou et al., 2023; Lee et al., 2024), we presented a set of storyboards to participants with the same background from which the design space was derived. Each storyboard depicted speculative scenarios combining different feedback modalities, body locations, exercise types, and sensing conditions. The goal was to validate and refine the design space with users, while surfacing the strengths and limitations of different design ideas across scenarios.
We developed ten storyboard scenarios (S1–S10) that instantiated different combinations of Sensing & Triggering and Feedback Delivery dimensions (see Fig. 11).
6.0.1. Study Protocol
We recruited seven beginners (B1–B7) with fewer than three months of consistent strength training experience to participate in three workshops (ages 24–34; 4 male, 3 female). Each workshop was conducted in our lab space, which was equipped with two pairs of dumbbells (2 and 12 pounds) and an adjustable workout bench. To establish context, participants were first asked to perform two exercises, dumbbell incline fly and lateral raise, by following short tutorial videos111https://www.youtube.com/@nutritioneering, as these movements cover two primary degrees of freedom of the shoulder. Participants were also free to try additional exercises from the tutorial set if desired.
Following this exercise phase, participants were presented with a series of storyboards illustrating scenarios from the FlexGuard design space. They were asked to reflect on whether the scenarios aligned with their needs, expectations, and perceived challenges in strength training. Participants could continue to access the equipment and tutorial videos throughout the session. Each workshop lasted approximately one hour, with the first 10 minutes dedicated to exercise practice and the remaining 50 minutes to storyboard discussion. Participants received $20 in compensation for their time.
6.1. Results
Two researchers analyzed the workshop transcripts. This analysis clarified and validated our assumptions about FlexGuard, allowing us to refine the proposed design space. In addition, we identified recurring themes that reflect underlying user needs and design considerations emerging from the speculative storyboard scenarios. These insights inform the design of future on-body feedback systems aimed at scaffolding safety during strength training, as participants reasoned about both becoming aware of safety-relevant movement and feeling physically supported during execution.
6.1.1. Semantics of Feedback Modalities
Participants suggested that different feedback modalities could carry distinct meanings beyond their mechanical effects. intrinsic mechanisms, such as those shown in S7, S8, and S9, were commonly perceived as intuitive cues for posture correction, as reflected in 5 out of 7 participants’ interpretations. As B2 noted when discussing S8, “If the material itself keeps me from going further, I immediately know what I’m doing wrong.”
In contrast, extrinsic cues were interpreted more symbolically. Vibration often conveyed that “something is wrong or needs attention,” prompting users to pause and reflect. For example, in bench press (S10), if the right hand drifted ahead of the left, vibration on the right arm could signal the user to pause and allow balance to be restored before continuing. Similarly, pressure applied to a muscle group in S3 was described as resembling a trainer’s tap to draw attention to the correct muscle.
Participants also speculated that inconsistent feedback semantics could cause confusion, as B7 remarked: “The feedback has to be consistent. Otherwise I won’t know what to do.” Several participants envisioned hybrid approaches that combine intrinsic mechanisms for corrective guidance with extrinsic cues for reminders or motivational prompts, highlighting the importance of consistent and interpretable semantics for scaffolding safety.
6.1.2. Movement Continuity
Participants frequently reasoned about on-body feedback in terms of whether it allowed them to move smoothly through an exercise or interrupted movement at critical moments. Rather than focusing on specific modalities or implementations, the feedback preferences were around maintaining movement continuity during safe execution and accepting interruption when safety risks increased (Figure 12).
During continuous movement, participants preferred feedback that minimally interfered with execution, allowing them to maintain rhythm and comfort while remaining aware of deviations. In these phases, lower levels of interference were favored, as participants felt able to adjust independently without disruption. For example, in S3, participants emphasized that feedback should remain subtle during execution: “It should be subtle so it doesn’t make me jump mid-movement, and once I’ve corrected it should go away.” Participants also suggested that feedback could relax or reset after the execution phase, reflecting a desire to avoid persistent interruption when movement was perceived as safe.
In contrast, when participants perceived heightened safety risk, such as fatigue-induced shaking or approaching range-of-motion boundaries, they valued feedback that interrupted movement continuity. In S5, participants appreciated feedback that provided physical assistance to help complete remaining repetitions while making limits explicit: “I’d like this feature—it’s similar to how peers spot each other” (B5). Similarly, in S8, participants described the usefulness of feedback that made further movement difficult at boundaries: “If I’m about to hurt myself, I want it to stop me right away.”
6.1.3. Adaptability Across Contexts
Participants’ reflections on feedback semantics and intensity suggested that adaptability is important across different contexts of use. For instance, when performing an exercise, participants preferred feedback that remains subtle during movement but becomes progressively stronger as deviations increase, favoring elastic feedback for flexibility (S7). At extrema positions, however, they preferred stronger constraints with rigid feedback to clearly signal safe ROM boundaries (S8). This indicates that participants imagined wearables that adapt not only across exercises, but also across different phases of movement.
Policies such as expert-based were mentioned as one way to guide such transitions. Participants imagined systems that could use thresholds for limiting ROM at both ends to escalate from elastic to rigid feedback once a boundary is reached (mentioned by B1, B2, B5, and B6). Participants also speculated about personalized policies, such as adapting feedback based on fatigue (S5) or changing reminders as the exercise and target region shift. Adaptability was also expected to vary with user expertise. For instance, one participant noted, “Once I’ve learned and memorized the correct movement, a reminder might be less necessary.”
6.1.4. Other Findings
Some scenarios, such as S2, were seen as useful in specific contexts. For example, when training with a mild injury, stabilization could help redistribute load away from affected muscles.
At the same time, participants reflected on how wearables might influence learning. Several expressed concern that excessive support could reduce opportunities to internalize correct form. For instance, in S2, participants noted a tension between learning through practice and appreciating stabilization that helps ensure correct muscle engagement. In contrast, S9 was regarded as more supportive of learning by providing feedback on tempo. As B1 noted, “I could know if I go too fast or not, because I think that’s important to know.”
Participants also raised concerns about potential risks associated with locking mechanisms such as in S6. If weights were falling and users were unable to reach a safe position while locked, this could be dangerous. As B5 explained, “I will hold myself accountable and not rush through the steps, so I might not use it.” To address this concern, B3 suggested including an “emergency button” that would release the lock when needed.
6.1.5. Summary
Together, these results validate the dimensions of the FlexGuard design space while also surfacing opportunities for refinement. Because the speed dating method relied on storyboards, participants’ reflections were necessarily speculative rather than grounded in embodied practice despite trying a few exercises in the lab. Nevertheless, adaptability emerged as a recurring priority, particularly for feedback that adjusts across phases of movement. The workshops also clarified the scope of what should be included within the design space. For example, we were initially uncertain whether feedback was necessary for inter-limb coordination (S10), since users can naturally perceive weight imbalance during an exercise. However, across all sessions, participants (B1, B2, B5, B7) consistently emphasized the value of on-body cues to confirm user’s interpretation of the situation—B5: “Although I can feel something is off, the vibration tells me exactly what’s happening.”
7. WORKSHOPS WITH A PROOF-OF-CONCEPT
We sought to translate FlexGuard into a proof-of-concept (PoC) prototype and utilize it during training to further validate how the design space operates in practice. In addition, we also aimed to gather additional evidence on users’ underlying needs, several of which were surfaced during the speed dating study. In this section, we outline the scope of this validation, the rationale behind the PoC design, and the study protocol and results.
7.1. Motivation for PoC
Given the breadth of FlexGuard, we focused our scope on Feedback Delivery, as it is the aspect that most requires direct user experience to evaluate in context. While prior work in physical training and related domains has examined portions of this design space, primarily through extrinsic feedback such as vibrotactile systems for snowboarding (Spelmezan et al., 2009) and rock climbing (Kosmalla et al., 2016) (e.g., extrinsic (discrete), instant or delayed alerts), high-fidelity full-body suits like the Teslasuit (Caserman et al., 2021) (e.g., continuous, instant cues), and pneumatic garments such as VabricBeads (Pardomuan et al., 2024) (e.g., extrinsic (continuous), compressive actuation), these systems have largely focused on delivering feedback rather than examining how feedback scaffolds trainees’ experience of safety during movement. While such systems deliver useful support, intrinsic feedback remains comparatively underexplored and yet offers distinct benefits for personalized, intelligent wearables that scaffold safety: it provides continuous, instant cues generated directly through the material properties of the wearable itself, allowing users to remain aware of safety-relevant movement while experiencing embodied support as part of their motion rather than as an external interruption.
Guided by these observations and insights from our previous study, we scoped the PoC to both validate design dimensions through an embodied instantiation of intrinsic feedback and clarify user-needs assumptions surfaced during speed dating. Because our focus is on Feedback Delivery, the PoC operationalizes a set of dimensions that extends beyond what prior work has addressed, including Feedback Type (intrinsic via both rigid and elastic mechanisms), Feedback Intensity (low, high, progressive), Feedback Latency (instant), and Feedback Direction (opposing, guiding), with Feedback Location at leading movement surfaces across the upper body.
7.2. PoC Suit Design
The PoC suit prototype consisted of a skin-fitting shirt equipped with rigid straps across the chest, shoulders, upper arms, and posterior forearms, complemented by elastic straps on the shoulders. The PoC suit was designed around lateral raise and incline dumbbell fly (Figure 13), as these exercises together span two key degrees of freedom of the shoulder, a region associated with common safety concerns and strain in strength training (Hedrick and Wada, 2008). To account for phase adaptation, each strap was mapped to different phases of movement. Although strap adjustments were performed manually by researchers, this setup simulated how a future motorized system could dynamically modulate intrinsic resistance in real time.
Rigid straps (r1–r4) functioned as opposing feedback, constraining movement in undesired directions and offering adjustable low and high intensity levels through strap tightening. In contrast, the elastic strap (l1) acted as guiding feedback, nudging the arm toward the intended posture with progressive intensity, where larger deviations resulted in stronger resistance.
For the incline dumbbell fly, r1 constrained elbow extension as the arms transitioned from folded to straightened positions, while r3 supported upright torso alignment during the hold phase. Strap r4 applied opposing feedback near the folding limit of the elbows, making excessive flexion difficult and indicating a boundary for elbow movement. In parallel, l1 offered guiding progressive resistance throughout the movement phases, helping participants remain aware of arm trajectory within the intended movement plane relative to the body.
In the lateral raise, r2 regulated arm elevation by linking the torso and shoulder, while r4 again constrained elbow flexion, limiting collapse and signaling acceptable movement ranges.
7.3. System Operation
The PoC system, as shown in Figure 14, operated through three phases: sensing, decision-making, and actuation. For sensing, two external cameras captured Front and Side views of participants’ movements, which were streamed to the PoC’s digital interface displayed on a monitor visible to the researchers. These video feeds were processed with MediaPipe (Dill et al., 2023) for real-time pose estimation. Using key postures for each exercise (see Supplementary Material), joint-angle comparisons were computed against ground-truth angles derived from tutorial videos, using the joint landmarks obtained from each camera view. When a real-time joint angle exceeded the allowable threshold in either view, the interface displayed the corresponding angle label in red; when the angle remained within bounds, the label appeared in green (Figure 14’s before-after).
Decision-making was driven by a system notification, illustrated in the notification panel in figure 14, computed from the magnitude and consistency of deviations from the ground-truth posture. For each set of five repetitions, the system calculated the deviation angle for every rep and considered a deviation meaningful only if it exceeded the threshold in a majority of repetitions (i.e., at least 3 out of 5). When this occurred, the system averaged the deviation angles across the above-threshold repetitions to determine the recommended feedback level for that joint. Small averaged deviations (e.g., –) triggered a recommendation for low-intensity feedback, whereas larger averaged deviations (e.g., or more) triggered high-intensity feedback. For motion-path deviations, the system applied an additional buffer to account for reduced MediaPipe performance in the Side view. As before, if more than half of the repetitions exceeded this adjusted threshold, the system recommended progressive feedback via the elastic strap. All of the information will be summarized in a notification after each set of exercise. When deviations did not meet the criteria, the system recommended no feedback for that joint.
We approximated the actuation phase by having researchers manually tighten the straps to simulate adaptive intrinsic feedback. Two researchers continuously monitored the interface, and after each set, the monitor displayed a notification indicating where strap adjustments were needed. Researchers then manually configured the feedback level on the relevant joint while the participant was idle in their rest posture between sets, as illustrated in the before and after comparison panel in Figure 14. We adopted this approach because fully automating intrinsic feedback actuation in real time, while keeping the system wearable and non-intrusive for strength training, is non-trivial and beyond the scope of this paper’s contribution.
7.4. Protocol
We recruited six beginners in strength training (fewer than three months of consistent training; 24M, 23M, 29M, 21F, 24F, 24M). The study was conducted in our lab, which was equipped with an adjustable workout bench and two sets of dumbbells. After providing informed consent, participants put on the PoC suit, with disposable elbow sleeves supplied to prevent direct skin contact. At least one researcher of the same gender as the participant assisted with putting on the suit and adjusting the straps. The session proceeded as follows:
Calibration: For each participant, we established baseline “low” and “high” feedback intensity levels and recorded their locations on the straps. We also measured rigid strap tension using a Newton spring scale.
Exercise Trials: Participants performed five repetitions of each exercise while following a tutorial video displayed on the monitor of the PoC platform. The PoC tracked the exercise phase-by-phase. When posture deviations occurred, researchers manually adjusted the straps to simulate adaptive feedback.
Repeated Trials: Participants repeated the exercise for another five repetitions at the adjusted intensity level to experience the adjusted feedback intensity.
Survey and Interview: After the PoC trial, participants completed a questionnaire on perceived support, comfort, and appropriateness of feedback. While completing the questionnaire, they were asked to think aloud, and we conducted follow-up questions to further explore their reflections on adaptability, effectiveness, and potential improvements. During this phase, participants were also free to retry exercises or explore feedback configurations on the suit that they had not experienced earlier in the session.
Each study session lasted approximately one hour, and participants received $20 in compensation for their time and effort.
7.5. Results
We analyzed the survey responses, and two researchers reviewed the interview recordings. This analysis focused on evaluating how the PoC validated the proposed design space dimensions outlined in the study scope. We also report recurring themes derived from the study data, some of which were previously identified in the speed dating study. More details about the PoC’s corrective effectiveness can be found from Supplementary Material.
7.5.1. Overview of Evaluated Dimensions
Each of the dimensions was validated through subjective data or objective data from the pose tracking system.
Feedback Type. The PoC evaluated two intrinsic mechanisms, rigid and elastic straps, to examine how participants perceived differences in constraint and flexibility during movement. Several participants described rigid feedback as a strong constraint that helped them avoid unsafe positions: one noted that the straps “helped restrict [me] to overdoing some position because the tension is kind of really tight and strong,” (P2), while another explained that they are “always afraid of overextending” and that the suit “definitely helped restrict me from doing that” (P3). At the same time, others criticized rigid straps as “too restrictive” or uncomfortable when bending the arm. Overall, 4 out of 6 participants (P1-P3, P5) described the rigid straps to limit unsafe motion. In contrast, participants characterized elastic feedback as more supportive than restrictive. One participant stated, “elastic is better because you can still get the feedback but it’s less… I don’t think I need to be restricted a ton, I just need a little bit of feedback and then I can adjust the movement myself,” and another preferred elastic straps because they “give you a pathway back to ground… if you pull, you feel that pull… from both angles” (P4). Overall, four participants expressed a preference for elastic configurations for at least some phases of movement.
Participants generally described rigid feedback as providing a “firm stop” that prevented unsafe motion (P2, P4), while elastic feedback offered “helpful guidance” that supported a natural trajectory without feeling restrictive (P3, P6).
Feedback Intensity. The PoC implemented three levels of Feedback Intensity, i.e., low, high, and progressive, to reflect varying degrees of corrective resistance. Participants described low intensity as a subtle prompt that let them adjust their own form, noting that they “just need a little bit of feedback” (P2) rather than being strongly restricted. high intensity was associated with strong stopping forces that “helped restrict [me] from overdoing some positions,” but several people also found the highest level “too restrictive” (P3) or “more uncomfortable” (P5) and asked either for more resistance in some cases or less in others. progressive intensity was experienced as feedback that increases gradually where one participant explained that with the bending strap “the feedback is from weak to strong,” and others said that they could feel the tension increase when they moved further out of range, which aligns with the survey data (Figure 15) that prefers adjustable or progressive resistance over a single fixed intensity level.
Feedback Latency. The system simulated instant feedback by manually tightening or relaxing straps immediately upon detecting form deviations. Participants highlighted the importance of timing in feedback delivery, describing immediate correction as essential to maintaining safe technique, P1 noted that the straps gave “strong feedback” that prevented them from “going the other way around” or “moving up over the final position,” and that they could “feel feedback appropriately aligned with the different phases of movement.” At the same time, several participants implicitly expressed preferences with delayed feedback, explaining that the suit often corrected posture at the boundaries of the motion rather than during the movement itself. For example, one participant explained, “it helped me fix my posture at the end,” (P2) and another remarked that they “didn’t feel that much feedback during my movements” (P6) and only noticed it “when I’m out of range.” These reflections indicate openness to delayed or boundary-based cues, especially when continuous correction felt unnecessary or overly restrictive.
Feedback Direction. Two directional strategies were tested through the PoC straps: opposing and guiding feedback. Five participants interpreted directional cues as meaningful signals to adjust their movement. opposing feedback provided resistance against incorrect motion paths, while guiding feedback assisted users back toward the correct trajectory. opposing feedback was noted by three participants as resistance that helped push them back if they drift off-path. P5 noted ”I can feel the tension if I’m in the wrong position” and P4 described it as ”You can feel the pull from both angles.” In contrast, guiding feedback was mentioned in four sessions, helping participants understand where their arms should remain without fully restrict the motion. ”It helped [me] restrict from overextending.” (P3), and P6 reported that the elastic strap ”the stretch helped me remain in the correct position instead of like just lying down.”
Together, these findings demonstrate that the PoC successfully embodied the adaptive principles of the FlexGuard design space, validating user needs for responsive, phase-aware, and semantically interpretable on-body feedback mechanisms in safety-oriented training.
7.5.2. Phase-Specific Adaptations
Participants strongly valued feedback that adapted to different phases of movement. Survey responses reflected this, with high ratings for “Feedback was appropriately aligned with different phases of movement” (mean = 5.7). As P1 confirmed, “I can properly feel it in the different phases as I move.” Similarly, P3 emphasized the importance of differential support: “When lowering the dumbbells, I need more resistance, but at the top I don’t want to feel held back.” Together, these accounts reinforce phase-specific adaptability as both perceptible and appreciated, establishing it as a core design principle.
7.5.3. Movement Continuity in Embodied Use
Consistent with the movement continuity considerations surfaced in the speed dating study, the PoC revealed a clear embodied tension between maintaining movement continuity and applying corrective constraint. The PoC was rated highly for corrective support (mean = 6.0) but lower for comfort (mean = 4.2). Higher levels of constraint (rigid feedback) were experienced as effective for restricting unsafe ranges of motion, but often at the cost of disrupting movement continuity. As P5 noted, “If it’s too strict, I feel like I can’t really move.”
In contrast, participants reported that elastic straps allowed them to continue moving while remaining aware of deviations, enabling adjustment without full interruption. However, this flexibility also introduced limitations. Several participants noted that elastic feedback could be too subtle near the intended plane of motion, contributing to a lower score on staying on the correct movement path (mean = 4.3). This effect was partly due to calibration constraints in the prototype. Despite this, participants still reported feeling guided during exercise (mean = 5.3), describing elastic feedback as providing a “kickback” force that nudged movement back on track rather than stopping it outright. The finding demonstrates how different feedback strategies shift the balance between allowing movement to continue and interrupting it when safety risks increase.
7.5.4. Scaffolding Safety-Relevant Movement
Participants recognized the system’s potential to scaffold safety by supporting their awareness of physical limits and providing embodied constraint at safety-critical moments. Participants rated highly on “The suit helped me avoid unsafe ranges of motion” (mean = 5.7). P2 explained, “This suit stops me before going too far.” Similarly, P4 emphasized the value of strong corrective cues: “The suit stops me right away, which is really great if I’m about to hurt myself.” Together, these results validate the role of rigid straps as movement constraints that scaffold trainees’ awareness of and response to safety-relevant movement at critical moments.
7.5.5. Suggestions and Improvements
While overall perceptions were positive, participants also identified areas for improvement. Comfort was a recurring concern, with several noting issues such as ill-fitting shirt sizes and strap tightness that could limit long-term use. As P3 remarked, “I like the idea, but I wouldn’t want to wear this for a whole workout—it gets uncomfortable. Maybe because the shirt is size M.” P6 also observed that when strap intensity increased, tension was transferred to anchoring points, causing discomfort in the forearm even on sides not directly involved. These reflections suggest opportunities for refining future prototypes to improve fit, comfort, and localized feedback delivery.
8. DISCUSSION
8.1. Implications for Sports and Fitness Wearables
Across our series of studies, we synthesized a set of design guidelines for future on-body feedback wearables that scaffold safety during strength training. These guidelines highlight how different forms of feedback function in practice and how they can be strategically combined to help users recognize safety-relevant movement while also providing embodied support during training.
8.1.1. Complementary Roles of Intrinsic and Extrinsic Feedback
Our findings show that intrinsic and extrinsic feedback modalities offer distinct strengths, challenges, and experiential qualities in how they scaffold safety during strength training. Table 3 summarizes these differences. These complementary qualities point toward hybrid systems that combine the immediacy and physical grounding of intrinsic feedback with the flexibility and interpretability of extrinsic cues to support trainees’ awareness of and response to safety-relevant movement.
| Dimension | Intrinsic Feedback | Extrinsic Feedback |
| User Experience | • Embodied guidance that stabilizes posture • Clear physical boundaries and ROM cues • May feel restrictive or uncomfortable at higher tensions | • Lightweight cues that do not restrict movement • Effective for timing, attention, and symbolic signaling • Risk of ambiguity if cue semantics are inconsistent |
| Technical Challenges | • Limited adjustability once configured due to reliance on material properties • Physical bulk or tension may impact comfort or mobility • More difficult to encode nuanced or varied feedback meanings | • Requires power, electronics, and careful placement • Harder to convey physical boundaries without relying on user interpretation • Timing and coordination must be carefully calibrated to avoid overload |
| Best-Suited Contexts | • Constraining ROM boundaries • Guiding slow or controlled movement • Providing stability during load-bearing tasks | • Signaling imbalance, tempo deviations, or fatigue • Supporting early-stage learning or technique awareness • Event-driven alerts and motivational or attentional cues |
8.1.2. User Needs for FlexGuard’s Application
While FlexGuard provides a structured way to reason about sensing and on-body feedback, applying the design space in practice requires attention to several user-centered needs. Below, we summarize the key considerations that emerged across our studies.
Adaptability to context. Wearable feedback should flexibly adjust to both rapid changes, such as transitions across movement phases, and slower-changing factors such as training experience, fatigue, or individual body mechanics. This adaptability may involve modulating feedback modality and intensity based on the moment-to-moment demands of the exercise. Subtle cues tended to be helpful during continuous motion, whereas stronger, more restrictive cues were more appropriate near ROM limits or under higher loads. These insights point toward context-aware systems that can modulate feedback in real time while still accommodating longer-term user characteristics.
Interpretable feedback semantics. Feedback should convey clear and consistent meaning to remain actionable. Some signals, such as rigid stops at ROM boundaries, naturally communicate their intent. However, physical constraints alone may not always indicate how to correct posture, guiding users closer to (but not precisely into) the intended position. Extrinsic modalities such as vibrotactile cues can express more nuanced information, though they also rely on how users interpret the signal. When users can understand what a cue means and how to adjust accordingly, feedback becomes a stronger support for learning safe movement patterns. Ensuring that feedback remains understandable is therefore an important design consideration, especially as guidance becomes more complex.
Movement autonomy. Effective systems should support freedom of movement when it is safe while offering firmer constraints when necessary. Excessive restriction during dynamic phases can feel uncomfortable or disrupt natural exercise rhythm, whereas carefully timed constraints at higher-risk points can help prevent unsafe ROM or posture breakdown. These observations suggest the value of systems that intervene selectively, preserving autonomy during safe motion while offering support when risk increases.
8.1.3. Toward Interactive Training Partner
By situating safety scaffolding within a design space framework, FlexGuard advances the conversation on how future sports and fitness wearables can evolve from passive tracking systems to interactive training partners. Our findings illustrate how wearables can move beyond performance monitoring to actively guide movement and support users in recognizing safety-relevant deviations during exercise, while still supporting motor learning. The idea has broader implications for adjacent domains such as rehabilitation, physical education, and recreational training, where embodied safety and skill acquisition are equally central. In this way, FlexGuard demonstrates how design spaces can serve as actionable frameworks for envisioning the next generation of intelligent, interactive wearables.
8.2. Intelligence in Materials
Our findings show that there is a need for greater intelligence in materials themselves. Many participant needs pointed toward adaptive intrinsic feedback, where feedback is conveyed through material properties such as rigidity or elasticity. Unlike extrinsic actuation, which can be modulated digitally or mechanically, intrinsic feedback relies on the physical response of the material. To make such feedback adaptive, materials themselves must become “smart” and dynamically tunable. This suggests a path forward through mechano- or physical intelligence (Alu et al., 2025), where material properties are engineered to support highly adaptive, wearable use cases.
8.3. Balancing Corrective and Supportive Feedback
There is a tension between feedback that is corrective and feedback that is supportive. Participants frequently contrasted rigid straps, which acted as “hard stops,” with elastic straps, which functioned more as guides. This mirrors pedagogical debates in coaching, where overcorrection may undermine motor learning while undercorrection risks unsafe execution. The implication for wearable design is the need for hybrid strategies that balance these roles, such as graduated resistance that escalates when errors persist or supportive cues that reinforce correct posture without being overbearing. Future sports and fitness wearables may need to embody this balance to ensure both safety and long-term skill acquisition.
9. LIMITATIONS AND FUTURE WORK
Our work carries several limitations that inform directions for future research. First, we investigated only upper-body movements. While these capture many common safety-sensitive patterns, they represent only a subset of the body. Some insights may generalize to other regions with similar roles, such as joints that govern ROM, supporting muscle groups, and related structures, but future work should extend our approach to lower-body and full-body exercises, where different biomechanical demands may surface new requirements for sensing and feedback design.
Second, in our PoC validation we relied on manual configuration of straps to set feedback intensity levels with six participants. While sufficient for this study, this approach introduced setup delays and required facilitator involvement. While we deliberately incorporated certain underexplored design dimensions into the PoC, there remains an opportunity for more comprehensive evaluation of other aspects, such as the semantics of various feedback modalities, especially in scenarios involving multiple simultaneous feedback cues. Coordinating such cues effectively to minimize cognitive load and maintain understandable semantics in time-sensitive settings represents an important next step toward supporting safe training. Additionally, higher system fidelity and a larger number of users may help move the vision closer to a usable product.
Third, our strap tension measurements (7–15 N across different locations) revealed an important limitation for wearable systems relying on passive, strap-based feedback. Stronger constraining cues require higher strap tensions, yet higher tensions make it more difficult to maintain comfort and stability without bulky motors or anchoring structures. Designing garments that can distribute forces evenly across the body while remaining flexible and wearable in dynamic exercise contexts remains an open challenge.
Future implementations should build on this foundation by automating configuration through adaptive calibration, onboard sensing, or machine learning to reduce setup burden and enable seamless, on-the-fly personalization. Such systems should remain lightweight and body-conforming to support strength training. This direction would move beyond design-space validation toward deployable wearable systems capable of delivering continuous, real-time guidance in everyday training settings.
10. CONCLUSION
This paper presented FlexGuard, a design space for on-body wearable feedback to scaffold safety in strength training, derived from nine co-design workshops with trainer–trainee pairs using a technology probe that explored both extrinsic and intrinsic modalities. FlexGuard is articulated through dimensions of Sensing & Triggering and Feedback Delivery, providing designers with a framework to relate safety-relevant movement risks to appropriate forms of on-body feedback. We conceptually validated this design space through a speed dating study with seven beginners, which highlighted how users reason about feedback semantics, movement continuity, and adaptation across fast-changing training contexts. We further examined selected dimensions through a proof-of-concept prototype evaluated with six beginners, where results substantiated the relevance of the proposed dimensions and clarified assumptions about users’ underlying needs surfaced in the speed dating study. Together, these findings demonstrate that FlexGuard offers a valid and useful framework for designing on-body feedback systems that scaffold trainees’ awareness of and support for safety-relevant movement during strength training.
Acknowledgements.
This work was supported by the Center for Human-Computer Interaction at Virginia Tech (CHCI@VT). We thank our participants for their valuable insights. We also thank Utkarsha Mohan, Xiaohang Tang, and Tong Wu for generously lending study equipment.References
- (1)
- Al Maimani and Roudaut (2017) Ahmed Al Maimani and Anne Roudaut. 2017. Frozen Suit: Designing a Changeable Stiffness Suit and its Application to Haptic Games. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 2440–2448. doi:10.1145/3025453.3025655
- Al-Sada et al. (2020) Mohammed Al-Sada, Keren Jiang, Shubhankar Ranade, Mohammed Kalkattawi, and Tatsuo Nakajima. 2020. HapticSnakes: multi-haptic feedback wearable robots for immersive virtual reality. Virtual Real. 24, 2 (June 2020), 191–209. doi:10.1007/s10055-019-00404-x
- Aliyana et al. (2024) Akshaya Kumar Aliyana, Danying Yang, Orathai Tangsirinaruenart, and George K. Stylios. 2024. A garment-integrated textile stitch-based strain sensor device, IoT-Enabled for enhanced wearable sportswear applications. Results in Engineering 23 (2024), 102794. doi:10.1016/j.rineng.2024.102794
- Alqarni (2019) Ahmed Mohammed Alqarni. 2019. Common injuries in resistance training. Saudi Journal of Sports Medicine 19, 2 (2019), 38–42.
- Alu et al. (2025) Andrea Alu, Andres F Arrieta, Emanuela Del Dottore, Michael D Dickey, Samuele Ferracin, Ryan L Harne, Helmut Hauser, Qiguang He, Jonathan Brigham Hopkins, Lance Hyatt, Suyi Li, Stefano Mariani, Barbara Mazzolai, Alessio Mondini, Aniket Pal, Daniel J Preston, Anoop Rajappan, Jordan R. Raney, Pedro Reis, Stephen Andrew Sarles, Metin Sitti, Uba K. Ubamanyu, Martin van Hecke, and Kon-Well Wang. 2025. Roadmap on embodying mechano-intelligence and computing in functional materials and structures. Smart Materials and Structures (2025). http://iopscience.iop.org/article/10.1088/1361-665X/adb7aa
- Amershi et al. (2014) Saleema Amershi, Maya Cakmak, W. Bradley Knox, and Todd Kulesza. 2014. Power to the People: The Role of Humans in Interactive Machine Learning. AI Magazine 35, 4 (2014), 105–120. arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1609/aimag.v35i4.2513 doi:10.1609/aimag.v35i4.2513
- Aslam et al. (2025) Sumaira Aslam, Jean De Dieu Habyarimana, and Shi Yong Bin. 2025. Neuromuscular adaptations to resistance training in elite versus recreational athletes. Frontiers in Physiology 16 (2025), 1598149.
- Barnes et al. (2021) Matthew J Barnes, Ashley Petterson, and Darryl J Cochrane. 2021. Peak power output and onset of muscle activation during high pull exercise. The Journal of Strength & Conditioning Research 35, 3 (2021), 675–679.
- Braun and Clarke (2019) Virginia Braun and Victoria Clarke. 2019. Reflecting on reflexive thematic analysis. Qualitative Research in Sport, Exercise and Health 11, 4 (2019), 589–597. arXiv:https://doi.org/10.1080/2159676X.2019.1628806 doi:10.1080/2159676X.2019.1628806
- Brooks et al. (2024) Jas Brooks, Alex Mazursky, Janice Hixon, and Pedro Lopes. 2024. Augmented Breathing via Thermal Feedback in the Nose. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 26, 11 pages. doi:10.1145/3654777.3676438
- Caserman et al. (2021) Polona Caserman, Clemens Krug, and Stefan Göbel. 2021. Recognizing full-body exercise execution errors using the teslasuit. Sensors 21, 24 (2021), 8389.
- Catapult (2025) Catapult. 2025. Capture, Analyze, Plan, Share every aspect of performance. Retrieved Jul 15, 2025 from https://www.catapult.com
- Chang et al. (2020) Zekun Chang, Tung D. Ta, Koya Narumi, Heeju Kim, Fuminori Okuya, Dongchi Li, Kunihiro Kato, Jie Qi, Yoshinobu Miyamoto, Kazuya Saito, and Yoshihiro Kawahara. 2020. Kirigami Haptic Swatches: Design Methods for Cut-and-Fold Haptic Feedback Mechanisms. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3313831.3376655
- Choi et al. (2022) Kyung Yun Choi, Neska ElHaouij, Jinmo Lee, Rosalind W. Picard, and Hiroshi Ishii. 2022. Design and Evaluation of a Clippable and Personalizable Pneumatic-haptic Feedback Device for Breathing Guidance. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 6, 1, Article 6 (March 2022), 36 pages. doi:10.1145/3517234
- Davidoff et al. (2007) Scott Davidoff, Min Kyung Lee, Anind K Dey, and John Zimmerman. 2007. Rapidly exploring application design through speed dating. In International conference on ubiquitous computing. Springer, 429–446.
- Delazio et al. (2018) Alexandra Delazio, Ken Nakagaki, Roberta L. Klatzky, Scott E. Hudson, Jill Fain Lehman, and Alanson P. Sample. 2018. Force Jacket: Pneumatically-Actuated Jacket for Embodied Haptic Experiences. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems (Montreal QC, Canada) (CHI ’18). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3173574.3173894
- Dill et al. (2023) Sebastian Dill, Andreas Rösch, Maurice Rohr, Gökhan Güney, Luisa De Witte, Elias Schwartz, and Christoph Hoog Antink. 2023. Accuracy evaluation of 3D pose estimation with mediapipe pose for physical exercises. In Current Directions in Biomedical Engineering, Vol. 9. De Gruyter, 563–566.
- Elsayed et al. (2020) Hesham Elsayed, Martin Weigel, Florian Müller, Martin Schmitz, Karola Marky, Sebastian Günther, Jan Riemann, and Max Mühlhäuser. 2020. VibroMap: Understanding the Spacing of Vibrotactile Actuators across the Body. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 4, 4, Article 125 (Dec. 2020), 16 pages. doi:10.1145/3432189
- Elvitigala et al. (2024) Don Samitha Elvitigala, Armağan Karahanoğlu, Andrii Matviienko, Laia Turmo Vidal, Dees Postma, Michael D Jones, Maria F. Montoya, Daniel Harrison, Lars Elbæk, Florian Daiber, Lisa Anneke Burr, Rakesh Patibanda, Paolo Buono, Perttu Hämäläinen, Robby Van Delden, Regina Bernhaupt, Xipei Ren, Vincent Van Rheden, Fabio Zambetta, Elise Van Den Hoven, Carine Lallemand, Dennis Reidsma, and Florian ‘Floyd’ Mueller. 2024. Grand Challenges in SportsHCI. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 312, 20 pages. doi:10.1145/3613904.3642050
- Elvitigala et al. (2019) Don Samitha Elvitigala, Denys J.C. Matthies, Löic David, Chamod Weerasinghe, and Suranga Nanayakkara. 2019. GymSoles: Improving Squats and Dead-Lifts by Visualizing the User’s Center of Pressure. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3290605.3300404
- Escudero (2024) Julio Andres Escudero. 2024. Haptic Scores: Non-Visual Ways of Experiencing Choreography. In Companion of the 2024 on ACM International Joint Conference on Pervasive and Ubiquitous Computing (Melbourne VIC, Australia) (UbiComp ’24). Association for Computing Machinery, New York, NY, USA, 281–285. doi:10.1145/3675094.3678368
- Esposito et al. (2018) Daniele Esposito, Emilio Andreozzi, Antonio Fratini, Gaetano D Gargiulo, Sergio Savino, Vincenzo Niola, and Paolo Bifulco. 2018. A Piezoresistive Sensor to Measure Muscle Contraction and Mechanomyography. Sensors 18, 8 (2018). doi:10.3390/s18082553
- Evke et al. (2019) Erin E Evke, Dilara Meli, and Max Shtein. 2019. Developable rotationally symmetric Kirigami-based structures as sensor platforms. Advanced Materials Technologies 4, 12 (2019), 1900563.
- Faigenbaum and Myer (2010) Avery D Faigenbaum and Gregory D Myer. 2010. Resistance training among young athletes: safety, efficacy and injury prevention effects. British journal of sports medicine 44, 1 (2010), 56–63.
- Fisher et al. (2023) James Fisher, Patroklos Androulakis-Korakakis, Jürgen Giessing, Eric Helms, Brad Schoenfeld, Dave Smith, and Richard Winett. 2023. Supervision during resistance training: A comparison of trainer and trainee perceptions. International Journal of Strength and Conditioning 3, 1 (2023).
- Fisher (2025) James P Fisher. 2025. Supervision During Strength Training—the Interplay with Facilitation, Feedback and Attentional Focus: A Narrative Review. Sports Medicine (2025), 1–9.
- Gabbett (2020) Tim J Gabbett. 2020. How much? How fast? How soon? Three simple concepts for progressing training loads to minimize injury risk and enhance performance. Journal of orthopaedic & sports physical therapy 50, 10 (2020), 570–573.
- Gillies (2019) Marco Gillies. 2019. Understanding the Role of Interactive Machine Learning in Movement Interaction Design. ACM Trans. Comput.-Hum. Interact. 26, 1, Article 5 (Feb. 2019), 34 pages. doi:10.1145/3287307
- Günther et al. (2019) Sebastian Günther, Mohit Makhija, Florian Müller, Dominik Schön, Max Mühlhäuser, and Markus Funk. 2019. PneumAct: Pneumatic Kinesthetic Actuation of Body Joints in Virtual Reality Environments. In Proceedings of the 2019 on Designing Interactive Systems Conference (San Diego, CA, USA) (DIS ’19). Association for Computing Machinery, New York, NY, USA, 227–240. doi:10.1145/3322276.3322302
- Hamdan et al. (2019) Nur Al-huda Hamdan, Adrian Wagner, Simon Voelker, Jürgen Steimle, and Jan Borchers. 2019. Springlets: Expressive, Flexible and Silent On-Skin Tactile Interfaces. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–14. doi:10.1145/3290605.3300718
- Han et al. (2023) Violet Yinuo Han, Abena Boadi-Agyemang, Yuyu Lin, David Lindlbauer, and Alexandra Ion. 2023. Parametric Haptics: Versatile Geometry-based Tactile Feedback Devices. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (San Francisco, CA, USA) (UIST ’23). Association for Computing Machinery, New York, NY, USA, Article 65, 13 pages. doi:10.1145/3586183.3606766
- Hartmann et al. (2007) Björn Hartmann, Leith Abdulla, Manas Mittal, and Scott R. Klemmer. 2007. Authoring sensor-based interactions by demonstration with direct manipulation and pattern recognition. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’07). Association for Computing Machinery, New York, NY, USA, 145–154. doi:10.1145/1240624.1240646
- Hassan et al. (2017) Mahmoud Hassan, Florian Daiber, Frederik Wiehr, Felix Kosmalla, and Antonio Krüger. 2017. FootStriker: An EMS-based Foot Strike Assistant for Running. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 1, 1, Article 2 (March 2017), 18 pages. doi:10.1145/3053332
- Hedrick and Wada (2008) Allen Hedrick and Hiroaki Wada. 2008. Weightlifting movements: do the benefits outweigh the risks? Strength & Conditioning Journal 30, 6 (2008), 26–35.
- Herbaut and Tuloup (2025) Alexis Herbaut and Edouard Tuloup. 2025. Effect of weightlifting belts on lumbar biomechanics and muscle activity in deadlift and squat. Sports Engineering 28, 1 (2025), 1–9.
- Höök et al. (2017) Kristina Höök, Caroline Hummels, Katherine Isbister, Patrizia Marti, Elena Márquez Segura, Martin Jonsson, Florian ’Floyd’ Mueller, Pedro A.N. Sanches, Thecla Schiphorst, Anna Ståhl, Dag Svanaes, Ambra Trotto, Marianne Graves Petersen, and Youn-kyung Lim. 2017. Soma-Based Design Theory. In Proceedings of the 2017 CHI Conference Extended Abstracts on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI EA ’17). Association for Computing Machinery, New York, NY, USA, 550–557. doi:10.1145/3027063.3027082
- Hu et al. (2025) Hongci Hu, Mengqi Jiang, Kai Lin, Kinor Shou-xiang Jiang, and Ziqian Bai. 2025. ReKnit-Care: A Seamless-Knitted Sensing Glove for Sensory Rehabilitation and Adaptive Haptic Feedback. In Proceedings of the Nineteenth International Conference on Tangible, Embedded, and Embodied Interaction (TEI ’25). Association for Computing Machinery, New York, NY, USA, Article 58, 7 pages. doi:10.1145/3689050.3705979
- Huang et al. (2025) Bingjian Huang, Siyi Ren, Yuewen Luo, Qilong Cheng, Hanfeng Cai, Yeqi Sang, Mauricio Sousa, Paul H Dietz, and Daniel Wigdor. 2025. VibraForge: A Scalable Prototyping Toolkit For Creating Spatialized Vibrotactile Feedback Systems. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 1137, 18 pages. doi:10.1145/3706598.3714273
- Hutchinson et al. (2003) Hilary Hutchinson, Wendy Mackay, Bo Westerlund, Benjamin B. Bederson, Allison Druin, Catherine Plaisant, Michel Beaudouin-Lafon, Stéphane Conversy, Helen Evans, Heiko Hansen, Nicolas Roussel, and Björn Eiderbäck. 2003. Technology probes: inspiring design for and with families. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Ft. Lauderdale, Florida, USA) (CHI ’03). Association for Computing Machinery, New York, NY, USA, 17–24. doi:10.1145/642611.642616
- Islam et al. (2022) Md Shafiqul Islam, Sang Won Lee, Samantha M. Harden, and Sol Lim. 2022. Effects of vibrotactile feedback on yoga practice. Frontiers in Sports and Active Living Volume 4 - 2022 (2022). doi:10.3389/fspor.2022.1005003
- Jo et al. (2023a) Hye-Young Jo, Laurenz Seidel, Michel Pahud, Mike Sinclair, and Andrea Bianchi. 2023a. FlowAR: How Different Augmented Reality Visualizations of Online Fitness Videos Support Flow for At-Home Yoga Exercises. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 469, 17 pages. doi:10.1145/3544548.3580897
- Jo et al. (2023b) Hye-Young Jo, Chan Hu Wie, Yejin Jang, Dong-Uk Kim, Yurim Son, and Yoonji Kim. 2023b. TrainerTap: Weightlifting Support System Prototype Simulating Personal Trainer’s Tactile and Auditory Guidance. In Adjunct Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (San Francisco, CA, USA) (UIST ’23 Adjunct). Association for Computing Machinery, New York, NY, USA, Article 14, 3 pages. doi:10.1145/3586182.3616644
- Johnson et al. (2023) Tate Johnson, Dinesh K Patel, Humphrey Yang, Umut Serdar Civici, Adriane Fernandes Minori, and Lining Yao. 2023. FlexTure: Designing Configurable and Dynamic Surface Features. In Proceedings of the 2023 ACM Designing Interactive Systems Conference (Pittsburgh, PA, USA) (DIS ’23). Association for Computing Machinery, New York, NY, USA, 580–593. doi:10.1145/3563657.3595995
- Kao et al. (2018) Hsin-Liu (Cindy) Kao, Miren Bamforth, David Kim, and Chris Schmandt. 2018. Skinmorph: texture-tunable on-skin interface through thin, programmable gel. In Proceedings of the 2018 ACM International Symposium on Wearable Computers (Singapore, Singapore) (ISWC ’18). Association for Computing Machinery, New York, NY, USA, 196–203. doi:10.1145/3267242.3267262
- Kaul and Rohs (2017) Oliver Beren Kaul and Michael Rohs. 2017. HapticHead: A Spherical Vibrotactile Grid around the Head for 3D Guidance in Virtual and Augmented Reality. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 3729–3740. doi:10.1145/3025453.3025684
- Keelawat and Suzuki (2024) Panayu Keelawat and Ryo Suzuki. 2024. Transforming Procedural Instructions into In-Situ Augmented Reality Guides with InstructAR. In Adjunct Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST Adjunct ’24). Association for Computing Machinery, New York, NY, USA, Article 70, 3 pages. doi:10.1145/3672539.3686321
- Kilic Afsar et al. (2021) Ozgun Kilic Afsar, Ali Shtarbanov, Hila Mor, Ken Nakagaki, Jack Forman, Karen Modrei, Seung Hee Jeong, Klas Hjort, Kristina Höök, and Hiroshi Ishii. 2021. OmniFiber: Integrated Fluidic Fiber Actuators for Weaving Movement based Interactions into the ‘Fabric of Everyday Life’. In The 34th Annual ACM Symposium on User Interface Software and Technology (Virtual Event, USA) (UIST ’21). Association for Computing Machinery, New York, NY, USA, 1010–1026. doi:10.1145/3472749.3474802
- Kim and Asbeck (2022) Hubert Kim and Alan T. Asbeck. 2022. The Effects of Torque Magnitude and Stiffness in Arm Guidance Through Joint Torque Feedback. IEEE Access 10 (2022), 5842–5854. doi:10.1109/ACCESS.2022.3141981
- Kim et al. (2023) Jiha Kim, Younho Nam, Jungeun Lee, Young-Joo Suh, and Inseok Hwang. 2023. ProxiFit: Proximity Magnetic Sensing Using a Single Commodity Mobile toward Holistic Weight Exercise Monitoring. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 3, Article 105 (Sept. 2023), 32 pages. doi:10.1145/3610920
- Knoblauch et al. (2014) Hubert Knoblauch, René T., and Bernt Schnettler. 2014. Video Analysis and Videography Qualitative Methods.
- Kosmalla et al. (2016) Felix Kosmalla, Frederik Wiehr, Florian Daiber, Antonio Krüger, and Markus Löchtefeld. 2016. ClimbAware: Investigating Perception and Acceptance of Wearables in Rock Climbing. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’16). Association for Computing Machinery, New York, NY, USA, 1097–1108. doi:10.1145/2858036.2858562
- Kowsar et al. (2019) Yousef Kowsar, Eduardo Velloso, Lars Kulik, and Christopher Leckie. 2019. LiftSmart: a monitoring and warning wearable for weight trainers. In Adjunct Proceedings of the 2019 ACM International Joint Conference on Pervasive and Ubiquitous Computing and Proceedings of the 2019 ACM International Symposium on Wearable Computers (London, United Kingdom) (UbiComp/ISWC ’19 Adjunct). Association for Computing Machinery, New York, NY, USA, 298–301. doi:10.1145/3341162.3343795
- Kwon et al. (2022) Junghan Kwon, Inrak Choi, Myungsun Park, Jeongin Moon, Bomin Jeong, Prabhat Pathak, Jooeun Ahn, and Yong-Lae Park. 2022. Selectively stiffening garments enabled by cellular composites. Advanced Materials Technologies 7, 9 (2022), 2101543.
- Latella et al. (2024) Christopher Latella, Joel Garrett, and Daniel van den Hoek. 2024. How technological impacts on performance have been managed in elite sport: a powerlifting example. Journal of Applied Physiology 137, 4 (2024), 821–822.
- Latella et al. (2019) Christopher Latella, Daniel Van Den Hoek, and Wei-Peng Teo. 2019. Differences in strength performance between novice and elite athletes: Evidence from powerlifters. The Journal of Strength & Conditioning Research 33 (2019), S103–S112.
- Lee et al. (2026) Hsuanling Lee, Jiakun Yu, Shurui Zheng, Te-Yan Wu, and Liang He. 2026. FluxLab: Creating 3D Printable Shape-Changing Devices with Integrated Deformation Sensing. In Proceedings of the Twentieth International Conference on Tangible, Embedded, and Embodied Interaction (Chicago, IL, USA) (TEI ’26). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3731459.3773331
- Lee et al. (2024) Soohwan Lee, Seoyeong Hwang, Ian Oakley, and Kyungho Lee. 2024. Expanding the Design Space of Vision-based Interactive Systems for Group Dance Practice. In Proceedings of the 2024 ACM Designing Interactive Systems Conference (Copenhagen, Denmark) (DIS ’24). Association for Computing Machinery, New York, NY, USA, 2768–2787. doi:10.1145/3643834.3661568
- Leiva et al. (2021) Germán Leiva, Jens Emil Grønbæk, Clemens Nylandsted Klokmose, Cuong Nguyen, Rubaiat Habib Kazi, and Paul Asente. 2021. Rapido: Prototyping Interactive AR Experiences through Programming by Demonstration. In The 34th Annual ACM Symposium on User Interface Software and Technology (Virtual Event, USA) (UIST ’21). Association for Computing Machinery, New York, NY, USA, 626–637. doi:10.1145/3472749.3474774
- Ley-Flores et al. (2024) Judith Ley-Flores, Laia Turmo Vidal, Elena Márquez Segura, Aneesha Singh, Frederic Bevilacqua, Francisco Cuadrado, Joaquín Roberto Díaz Durán, Omar Valdiviezo-Hernández, Milagrosa Sánchez-Martin, and Ana Tajadura-Jiménez. 2024. Co-Designing Sensory Feedback for Wearables to Support Physical Activity through Body Sensations. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 1, Article 40 (March 2024), 31 pages. doi:10.1145/3643499
- Li et al. (2017) Shuguang Li, Daniel M. Vogt, Daniela Rus, and Robert J. Wood. 2017. Fluid-driven origami-inspired artificial muscles. Proceedings of the National Academy of Sciences 114, 50 (2017), 13132–13137. arXiv:https://www.pnas.org/doi/pdf/10.1073/pnas.1713450114 doi:10.1073/pnas.1713450114
- Li and Luo (2024) Yunzhu Li and Yiyue Luo. 2024. Intelligent textiles are looking bright. Science 384, 6691 (2024), 29–30. arXiv:https://www.science.org/doi/pdf/10.1126/science.ado5922 doi:10.1126/science.ado5922
- Lin et al. (2024a) Yuyu Lin, Jesse T Gonzalez, Zhitong Cui, Yash Rajeev Banka, and Alexandra Ion. 2024a. ConeAct: A Multistable Actuator for Dynamic Materials. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 324, 16 pages. doi:10.1145/3613904.3642949
- Lin et al. (2025a) Yuyu Lin, Hatice Gokcen Guner, Jianzhe Gu, Sonia Prashant, and Alexandra Ion. 2025a. Wearable Material Properties: Passive Wearable Microstructures as Adaptable Interfaces for the Physical Environment. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 305, 16 pages. doi:10.1145/3706598.3714215
- Lin et al. (2024b) Yilong Lin, Peng Zhang, Eyal Ofek, and Seungwoo Je. 2024b. ArmDeformation: Inducing the Sensation of Arm Deformation in Virtual Reality Using Skin-Stretching. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 407, 18 pages. doi:10.1145/3613904.3642518
- Lin et al. (2025b) Yuyu Lin, Dian Zhu, Anoushka Naidu, Kenneth Yu, Deon Harper, Eni Halilaj, Douglas Weber, Deborah Ellen Kenney, Adam J. Popchak, Mark Baratz, and Alexandra Ion. 2025b. Personalized Bistable Orthoses for Rehabilitation of Finger Joints. In Proceedings of the 38th Annual ACM Symposium on User Interface Software and Technology (Busan, Republic of Korea) (UIST ’25). Association for Computing Machinery, New York, NY, USA. doi:10.1145/3746059.3747643
- Lu et al. (2019) Qiuyu Lu, Jifei Ou, João Wilbert, André Haben, Haipeng Mi, and Hiroshi Ishii. 2019. milliMorph – Fluid-Driven Thin Film Shape-Change Materials for Interaction Design. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology (New Orleans, LA, USA) (UIST ’19). Association for Computing Machinery, New York, NY, USA, 663–672. doi:10.1145/3332165.3347956
- Luo et al. (2024) Yiyue Luo, Chao Liu, Young Joong Lee, Joseph DelPreto, Kui Wu, Michael Foshey, Daniela Rus, Tomás Palacios, Yunzhu Li, Antonio Torralba, et al. 2024. Adaptive tactile interaction transfer via digitally embroidered smart gloves. Nature communications 15, 1 (2024), 868.
- Ma et al. (2024) Dizhi Ma, Xiyun Hu, Jingyu Shi, Mayank Patel, Rahul Jain, Ziyi Liu, Zhengzhe Zhu, and Karthik Ramani. 2024. avaTTAR: Table Tennis Stroke Training with Embodied and Detached Visualization in Augmented Reality. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 35, 16 pages. doi:10.1145/3654777.3676400
- Machek et al. (2021) Steven B Machek, Thomas D Cardaci, Dylan T Wilburn, Mitchell C Cholewinski, Scarlett Lin Latt, Dillon R Harris, and Darryn S Willoughby. 2021. Neoprene knee sleeves of varying tightness augment barbell squat one repetition maximum performance without improving other indices of muscular strength, power, or endurance. The Journal of Strength & Conditioning Research 35 (2021), S6–S15.
- Mahadevan et al. (2022) Karthik Mahadevan, Yan Chen, Maya Cakmak, Anthony Tang, and Tovi Grossman. 2022. Mimic: In-Situ Recording and Re-Use of Demonstrations to Support Robot Teleoperation. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology (Bend, OR, USA) (UIST ’22). Association for Computing Machinery, New York, NY, USA, Article 40, 13 pages. doi:10.1145/3526113.3545639
- Mahadevan et al. (2024) Karthik Mahadevan, Jonathan Chien, Noah Brown, Zhuo Xu, Carolina Parada, Fei Xia, Andy Zeng, Leila Takayama, and Dorsa Sadigh. 2024. Generative Expressive Robot Behaviors using Large Language Models. In Proceedings of the 2024 ACM/IEEE International Conference on Human-Robot Interaction (Boulder, CO, USA) (HRI ’24). Association for Computing Machinery, New York, NY, USA, 482–491. doi:10.1145/3610977.3634999
- Morris et al. (2014) Dan Morris, T. Scott Saponas, Andrew Guillory, and Ilya Kelner. 2014. RecoFit: using a wearable sensor to find, recognize, and count repetitive exercises. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Toronto, Ontario, Canada) (CHI ’14). Association for Computing Machinery, New York, NY, USA, 3225–3234. doi:10.1145/2556288.2557116
- Paay et al. (2020) Jeni Paay, Jesper Kjeldskov, Frederick Sorensen, Thomas Jensen, and Oren Tirosh. 2020. Weight-Mate: Adaptive Training Support for Weight Lifting. In Proceedings of the 31st Australian Conference on Human-Computer-Interaction (Fremantle, WA, Australia) (OzCHI ’19). Association for Computing Machinery, New York, NY, USA, 95–105. doi:10.1145/3369457.3369466
- Pardomuan et al. (2024) Jefferson Pardomuan, Shio Miyafuji, Nobuhiro Takahashi, and Hideki Koike. 2024. VabricBeads: Variable Stiffness Structured Fabric using Artificial Muscle in Woven Beads. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 335, 17 pages. doi:10.1145/3613904.3642401
- Park et al. (2020) Junghoon Park, Junhwan Choi, Sangjoon J. Kim, Kap-Ho Seo, and Jung Kim. 2020. Design of an Inflatable Wrinkle Actuator With Fast Inflation/Deflation Responses for Wearable Suits. IEEE Robotics and Automation Letters 5, 3 (2020), 3799–3805. doi:10.1109/LRA.2020.2976299
- Payne et al. (2021) William Christopher Payne, Yoav Bergner, Mary Etta West, Carlie Charp, R. Benjamin Shapiro, Danielle Albers Szafir, Edd V. Taylor, and Kayla DesPortes. 2021. danceON: Culturally Responsive Creative Computing. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 96, 16 pages. doi:10.1145/3411764.3445149
- Peiris et al. (2017) Roshan Lalintha Peiris, Wei Peng, Zikun Chen, Liwei Chan, and Kouta Minamizawa. 2017. ThermoVR: Exploring Integrated Thermal Haptic Feedback with Head Mounted Displays. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17). Association for Computing Machinery, New York, NY, USA, 5452–5456. doi:10.1145/3025453.3025824
- Pu et al. (2025) Kevin Pu, Daniel Lazaro, Ian Arawjo, Haijun Xia, Ziang Xiao, Tovi Grossman, and Yan Chen. 2025. Assistance or Disruption? Exploring and Evaluating the Design and Trade-offs of Proactive AI Programming Support. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 152, 21 pages. doi:10.1145/3706598.3713357
- Rao et al. (2024) SK Rao, R Savalgi, VK Suresh, N Pyati, and S Krishnamurthy. 2024. FormFit: A physical exercise form correction system using BlazePose. In Computer Science Engineering. CRC Press, 490–497.
- Rudolph et al. (2022) Julius Cosmo Romeo Rudolph, David Holman, Bruno De Araujo, Ricardo Jota, Daniel Wigdor, and Valkyrie Savage. 2022. Sensing Hand Interactions with Everyday Objects by Profiling Wrist Topography. In Proceedings of the Sixteenth International Conference on Tangible, Embedded, and Embodied Interaction (Daejeon, Republic of Korea) (TEI ’22). Association for Computing Machinery, New York, NY, USA, Article 14, 14 pages. doi:10.1145/3490149.3501320
- Saini et al. (2024) Aryan Saini, Rakesh Patibanda, Nathalie Overdevest, Elise Van Den Hoven, and Florian ‘Floyd’ Mueller. 2024. PneuMa: Designing Pneumatic Bodily Extensions for Supporting Movement in Everyday Life. In Proceedings of the Eighteenth International Conference on Tangible, Embedded, and Embodied Interaction (Cork, Ireland) (TEI ’24). Association for Computing Machinery, New York, NY, USA, Article 1, 16 pages. doi:10.1145/3623509.3633349
- Seçkin et al. (2023) Ahmet Çağdaş Seçkin, Bahar Ateş, and Mine Seçkin. 2023. Review on Wearable Technology in sports: Concepts, Challenges and opportunities. Applied sciences 13, 18 (2023), 10399.
- Semeraro and Turmo Vidal (2022) Alessandra Semeraro and Laia Turmo Vidal. 2022. Visualizing Instructions for Physical Training: Exploring Visual Cues to Support Movement Learning from Instructional Videos. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 90, 16 pages. doi:10.1145/3491102.3517735
- Snouzy (2025) Snouzy. 2025. Modern open‐source fitness coaching platform. Create workout plans, track progress, and access a comprehensive exercise database. https://www.workout.cool/. Accessed: 2025-09-04.
- Spelmezan et al. (2009) Daniel Spelmezan, Mareike Jacobs, Anke Hilgers, and Jan Borchers. 2009. Tactile motion instructions for physical activities. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Boston, MA, USA) (CHI ’09). Association for Computing Machinery, New York, NY, USA, 2243–2252. doi:10.1145/1518701.1519044
- Stawarz et al. (2025) Katarzyna Stawarz, Aluna Everitt, and Judy Bowen. 2025. Exploring Opportunities for Flexible Wearables to Support Physical Training. In Proceedings of the 2025 ACM Designing Interactive Systems Conference (DIS ’25). Association for Computing Machinery, New York, NY, USA, 2156–2170. doi:10.1145/3715336.3735703
- Sung et al. (2024) Youjin Sung, Rachel Kim, Kun Woo Song, Yitian Shao, and Sang Ho Yoon. 2024. HapticPilot: Authoring In-situ Hand Posture-Adaptive Vibrotactile Feedback for Virtual Reality. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 4, Article 179 (Jan. 2024), 28 pages. doi:10.1145/3631453
- Tajadura-Jimenez et al. (2022) Ana Tajadura-Jimenez, Judith Ley-Flores, Omar Valdiviezo, Aneesha Singh, Milagrosa Sanchez-Martin, Joaquin Diaz Duran, and Elena Márquez Segura. 2022. Exploring the Design Space for Body Transformation Wearables to Support Physical Activity through Sensitizing and Bodystorming. In Proceedings of the 8th International Conference on Movement and Computing (Chicago, IL, USA) (MOCO ’22). Association for Computing Machinery, New York, NY, USA, Article 23, 9 pages. doi:10.1145/3537972.3538001
- Takahashi et al. (2024) Akifumi Takahashi, Yudai Tanaka, Archit Tamhane, Alan Shen, Shan-Yuan Teng, and Pedro Lopes. 2024. Can a Smartwatch Move Your Fingers? Compact and Practical Electrical Muscle Stimulation in a Smartwatch. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 2, 15 pages. doi:10.1145/3654777.3676373
- Tennent et al. (2020) Paul Tennent, Joe Marshall, Vasiliki Tsaknaki, Charles Windlin, Kristina Höök, and Miquel Alfaras. 2020. Soma Design and Sensory Misalignment. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–12. doi:10.1145/3313831.3376812
- Tessmer et al. (2022) Lavender Tessmer, Ganit Goldstein, Guillermo Herrera-Arcos, Volodymyr Korolovych, Rachel Bellisle, Cody Paige, Christopher Shallal, Atharva Sahasrabudhe, Hugh Herr, Svetlana Boriskina, Dava Newman, and Skylar Tibbits. 2022. 3D Knit Spacesuit Sleeve with Multifunctional Fibers and Tunable Compression.
- Tsai et al. (2022) Hsin-Ruey Tsai, Yu-So Liao, and Chieh Tsai. 2022. ImpactVest: Rendering Spatio-Temporal Multilevel Impact Force Feedback on Body in VR. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 356, 11 pages. doi:10.1145/3491102.3501971
- Turmo Vidal et al. (2023) Laia Turmo Vidal, Yinchu Li, Martin Stojanov, Karin B Johansson, Beatrice Tylstedt, and Lina Eklund. 2023. Towards Advancing Body Maps as Research Tool for Interaction Design. In Proceedings of the Seventeenth International Conference on Tangible, Embedded, and Embodied Interaction (Warsaw, Poland) (TEI ’23). Association for Computing Machinery, New York, NY, USA, Article 20, 14 pages. doi:10.1145/3569009.3573838
- Turmo Vidal et al. (2025) Laia Turmo Vidal, Annika Waern, Rosa Cabanas-Valdés, Lauren van Loo, Yinchu Li, and Karthik Venkataraman Meenaakshisundaram. 2025. Towards Personalized Physiotherapy through Interactive Machine Learning: A Conceptual Infrastructure Design for In-Clinic and Out-of-Clinic Support. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 670, 19 pages. doi:10.1145/3706598.3713823
- Turmo Vidal et al. (2020) Laia Turmo Vidal, Hui Zhu, and Abraham Riego-Delgado. 2020. BodyLights: Open-Ended Augmented Feedback to Support Training Towards a Correct Exercise Execution. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. doi:10.1145/3313831.3376268
- Turmo Vidal et al. (2021) Laia Turmo Vidal, Hui Zhu, Annika Waern, and Elena Márquez Segura. 2021. The Design Space of Wearables for Sports and Fitness Practices. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21). Association for Computing Machinery, New York, NY, USA, Article 267, 14 pages. doi:10.1145/3411764.3445700
- Velloso et al. (2013) Eduardo Velloso, Andreas Bulling, and Hans Gellersen. 2013. MotionMA: motion modelling and analysis by demonstration. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Paris, France) (CHI ’13). Association for Computing Machinery, New York, NY, USA, 1309–1318. doi:10.1145/2470654.2466171
- Wang et al. (2020) Chi Wang, Da-Yuan Huang, Shuo-Wen Hsu, Cheng-Lung Lin, Yeu-Luen Chiu, Chu-En Hou, and Bing-Yu Chen. 2020. Gaiters: Exploring Skin Stretch Feedback on Legs for Enhancing Virtual Reality Experiences. In Proceedings of the 2020 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’20). Association for Computing Machinery, New York, NY, USA, 1–14. doi:10.1145/3313831.3376865
- Wang et al. (2023) Guanyun Wang, Yue Yang, Mengyan Guo, Kuangqi Zhu, Zihan Yan, Qiang Cui, Zihong Zhou, Junzhe Ji, Jiaji Li, Danli Luo, Deying Pan, Yitao Fan, Teng Han, Ye Tao, and Lingyun Sun. 2023. ThermoFit: Thermoforming Smart Orthoses via Metamaterial Structures for Body-Fitting and Component-Adjusting. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 1, Article 31 (March 2023), 27 pages. doi:10.1145/3580806
- Wang et al. (2016) Qi Wang, Marina Toeters, Wei Chen, Annick Timmermans, and Panos Markopoulos. 2016. Zishi: A Smart Garment for Posture Monitoring. In Proceedings of the 2016 CHI Conference Extended Abstracts on Human Factors in Computing Systems (San Jose, California, USA) (CHI EA ’16). Association for Computing Machinery, New York, NY, USA, 3792–3795. doi:10.1145/2851581.2890262
- Wang et al. (2025) Zeyu Wang, Ruotong Yu, Xiangyang Wang, Jiexin Ding, Jiankai Tang, Jun Fang, Zhe He, Zhuojun Li, Tobias Röddiger, Weiye Xu, Xiyuxing Zhang, Huan-ang Gao, Nan Gao, Chun Yu, Yuanchun Shi, and Yuntao Wang. 2025. Computing with Smart Rings: A Systematic Literature Review. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 9, 3, Article 137 (Sept. 2025), 54 pages. doi:10.1145/3749480
- Weng et al. (2025) Jian-Jia Weng, Calvin Ku, Jo Chien Wang, Chih-Jen Cheng, Tica Lin, Yu-An Su, Tsung-Hsun Tsai, You-Yi Lin, Lun-Wei Ku, Hung-Kuo Chu, and Min-Chun Hu. 2025. Bridging Coaching Knowledge and AI Feedback to Enhance Motor Learning in Basketball Shooting Mechanics Through a Knowledge-Based SOP Framework. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 989, 20 pages. doi:10.1145/3706598.3713324
- Windlin et al. (2022) Charles Windlin, Kristina Höök, and Jarmo Laaksolahti. 2022. SKETCHING SOMA BITS. In Proceedings of the 2022 ACM Designing Interactive Systems Conference (Virtual Event, Australia) (DIS ’22). Association for Computing Machinery, New York, NY, USA, 1758–1772. doi:10.1145/3532106.3533510
- Wittchen et al. (2025) Dennis Wittchen, Alexander Ramian, Nihar Sabnis, Richard Böhme, Christopher Chlebowski, Georg Freitag, Bruno Fruchard, and Donald Degraen. 2025. CollabJam: Studying Collaborative Haptic Experience Design for On-Body Vibrotactile Patterns. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 1132, 20 pages. doi:10.1145/3706598.3713469
- Woźniak et al. (2020) Mikołaj P. Woźniak, Julia Dominiak, Michał Pieprzowski, Piotr Ładoński, Krzysztof Grudzień, Lars Lischke, Andrzej Romanowski, and Paweł W. Woźniak. 2020. Subtletee: Augmenting Posture Awareness for Beginner Golfers. Proc. ACM Hum.-Comput. Interact. 4, ISS, Article 204 (Nov. 2020), 24 pages. doi:10.1145/3427332
- Wu et al. (2024) Di Wu, Emily Guan, Yunjia Zhang, Hsuanju Lai, Qiuyu Lu, and Lining Yao. 2024. Waxpaper Actuator: Sequentially and Conditionally Programmable Wax Paper for Morphing Interfaces. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 865, 16 pages. doi:10.1145/3613904.3642373
- Xu et al. (2025) Ziao Xu, Chentian Zhang, Faqiang Wang, Jianyong Yu, Gang Yang, Roman A Surmenev, Zhaoling Li, and Bin Ding. 2025. Smart textiles for personalized sports and healthcare. Nano-Micro Letters 17, 1 (2025), 1–39.
- Yang et al. (2024) Yue Yang, Lei Ren, Chuang Chen, Bin Hu, Zhuoyi Zhang, Xinyan Li, Yanchen Shen, Kuangqi Zhu, Junzhe Ji, Yuyang Zhang, Yongbo Ni, Jiayi Wu, Qi Wang, Jiang Wu, Lingyun Sun, Ye Tao, and Guanyun Wang. 2024. SnapInflatables: Designing Inflatables with Snap-through Instability for Responsive Interaction. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 342, 15 pages. doi:10.1145/3613904.3642933
- Yang et al. (2023) Yue Yang, Lei Ren, Chuang Chen, Xinyue Wang, Yitao Fan, Yilin Shao, Kuangqi Zhu, Jiaji Li, Qi Wang, Lingyun Sun, Ye Tao, and Guanyun Wang. 2023. E-Orthosis: Augmenting Off-the-Shelf Orthoses with Electronics. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 625, 15 pages. doi:10.1145/3544548.3581471
- Yu et al. (2025) Tianhong Catherine Yu, Manru Mary Zhang, Luis Miguel Malenab, Chi-Jung Lee, Jacky Hao Jiang, Ruidong Zhang, François Guimbretière, and Cheng Zhang. 2025. SeamFit: Towards Practical Smart Clothing for Automatic Exercise Logging. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 9, 1, Article 24 (March 2025), 22 pages. doi:10.1145/3712287
- Yu et al. (2024) Xingyao Yu, Benjamin Lee, and Michael Sedlmair. 2024. Design Space of Visual Feedforward And Corrective Feedback in XR-Based Motion Guidance Systems. In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24). Association for Computing Machinery, New York, NY, USA, Article 723, 15 pages. doi:10.1145/3613904.3642143
- Zhang and Sra (2021) Bowen Zhang and Misha Sra. 2021. PneuMod: A Modular Haptic Device with Localized Pressure and Thermal Feedback. In Proceedings of the 27th ACM Symposium on Virtual Reality Software and Technology (Osaka, Japan) (VRST ’21). Association for Computing Machinery, New York, NY, USA, Article 30, 7 pages. doi:10.1145/3489849.3489857
- Zhang et al. (2024) Zining Zhang, Jiasheng Li, Zeyu Yan, Jun Nishida, and Huaishu Peng. 2024. JetUnit: Rendering Diverse Force Feedback in Virtual Reality Using Water Jets. In Proceedings of the 37th Annual ACM Symposium on User Interface Software and Technology (Pittsburgh, PA, USA) (UIST ’24). Association for Computing Machinery, New York, NY, USA, Article 136, 15 pages. doi:10.1145/3654777.3676440
- Zhou et al. (2023) Qiushi Zhou, Louise Grebel, Andrew Irlitti, Julie Ann Minaai, Jorge Goncalves, and Eduardo Velloso. 2023. Here and Now: Creating Improvisational Dance Movements with a Mixed Reality Mirror. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 183, 16 pages. doi:10.1145/3544548.3580666
- Zhou et al. (2025) Ran Zhou, Jianru Ding, Chenfeng Gao, Wanli Qian, Benjamin Erickson, Madeline Balaam, Daniel Leithinger, and Ken Nakagaki. 2025. Shape-Kit: A Design Toolkit for Crafting On-Body Expressive Haptics. In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (CHI ’25). Association for Computing Machinery, New York, NY, USA, Article 504, 26 pages. doi:10.1145/3706598.3713981
- Zhu et al. (2022) Junyi Zhu, Yuxuan Lei, Aashini Shah, Gila Schein, Hamid Ghaednia, Joseph Schwab, Casper Harteveld, and Stefanie Mueller. 2022. MuscleRehab: Improving Unsupervised Physical Rehabilitation by Monitoring and Visualizing Muscle Engagement. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology (Bend, OR, USA) (UIST ’22). Association for Computing Machinery, New York, NY, USA, Article 33, 14 pages. doi:10.1145/3526113.3545705